### v2

# Exploring Prompting Methods for Mitigating Class Imbalance through Synthetic Data Generation with Large Language Models

Abstract
Large language models (LLMs) have demonstrated impressive in-context learning capabilities across various domains. Inspired by this, our study explores the effectiveness of LLMs in generating realistic tabular data to mitigate class imbalance. We investigate and identify key prompt design elements such as data format, class presentation, and variable mapping to optimize the generation performance. Our findings indicate that using CSV format, balancing classes, and employing unique variable mapping produces realistic and reliable data, significantly enhancing machine learning performance for minor classes in imbalanced datasets. Additionally, these approaches improve the stability and efficiency of LLM data generation. We validate our approach using six real-world datasets and a toy dataset, achieving state-of-the-art performance in classification tasks. The code is available here.
Figure 1: Generation results on an imbalanced toy dataset. The major class is shown in blue, and the minor class in red. Our approach via in-context learning of LLM achieves (1) clear class boundaries, (2) accurate feature correlations, (3) proper value ranges, (4) well-represented relationships between categorical and numerical variables (last row), and (5) better diversity compared to the fine-tuned GReaT [4] model. Detailed explanations and complete results are provided in Appendix E.
1 Introduction
Tabular data plays a crucial role in various classification tasks, such as disease diagnosis [1] and financial analysis [3, 24]. However, challenges such as data scarcity and class imbalance, particularly for rare yet crucial events, can severely degrade the performance of machine learning (ML) models. To address these, synthetic data generation methods have been developed to augment tabular datasets cost-effectively. Traditional methods such as SMOTE and its variants [6, 12, 20] focus on generating data for minority classes to alleviate class imbalance. Recently, advanced deep learning-based generation models have shown promising results, such as TVAE [30], CTAB-GAN [33], and TabDDPM [18].
Large language models (LLMs) have also demonstrated significant potential in generating realistic tabular data, showing robust performance across numerical and categorical data [4, 31]. However, these models often require extensive, data-specific fine-tuning, which can lead to overfitting on major classes or dominant feature values when applied to small and imbalanced tabular datasets. An alternative approach leverages the in-context learning capabilities of LLMs. Extensive studies have demonstrated that LLMs can solve complex problems, learn and mimic patterns from input data, and generate new data without parameter updates [13, 17, 19, 28, 29, 32]. This capability holds significant potential for generating realistic and reliable tabular data to address class imbalance.
However, the in-context learning capabilities of LLMs largely depend on a prompt design [10, 26, 28, 32], while these methods are not directly applicable to tabular data generation, especially for imbalanced class scenarios. This is because tabular data is not inherently in plain text and involves unique challenges such as identifying feature correlations and accurately representing underrepresented attributes. Therefore, carefully crafted prompt methods specifically for imbalanced tabular data are essential to maximize LLM capabilities in generating realistic data effectively. While several studies have highlighted the potential of LLMs for generating synthetic tabular data using in-context learning [25, 31], they have not sufficiently explored various prompt design elements, such as data format or class presentation. This study aims to investigate the effectiveness of LLMs in resolving class imbalance problems and identify the optimal prompts to maximize this effectiveness.
Figure 2: Comparison of ML classification performance when adding synthetic data to the Travel dataset. The results are averaged across four classifiers, XGBoost, CatBoost, LightGBM, and Gradient Boosting Classifier. Each classifier is run five times, resulting in a total of 20 experiments.
To answer these questions, we explore various prompt methods and analyze the generated results to identify specific prompt structures and templates that significantly improve the quality of synthetic tabular data generated by LLMs, thereby enhancing ML classification performance. We propose several prompt design elements, including data format, class presentation methods, variable mapping, and task specification. Our findings indicate that providing data in CSV format, balancing and grouping classes, and applying a unique variable mapper to distinguish different variables significantly improve the quality and efficiency of synthetic data generation, underscoring their effectiveness in optimizing LLM performance. As shown in Fig. 1, the data generated using this prompt design exhibit the most distinct boundaries between target classes, accurate feature correlation, and variable ranges.
Moreover, this approach significantly improves the classification performance of ML models when applied to real-world datasets. As shown in Fig. 2, baseline methods exhibit low sensitivity, indicating that they fail to accurately generate minority class samples due to inherent class biases. In contrast, our method overcomes these, achieving high sensitivity and balanced performance for all classes.
In summary, our contributions are:
  • Our study explores and analyzes the effectiveness of LLMs in generating realistic synthetic tabular data via in-context learning, offering guidelines for designing prompts tailored to effectively exploit LLMs for this task, particularly addressing class imbalance.
  • We propose simple yet effective prompting methods that generate high-quality class-specific tabular data, maintaining feature correlation and value ranges.
  • The proposed approach significantly improves ML classification performance and enhances data generation efficiency and stability via in-context learning.
  • The proposed approach is generally applicable and easy to implement for any tabular classification task with minimal pre-processing, providing publicly available code.
  • We validate the effectiveness of our approach through extensive evaluation across six real-world public tabular data classification datasets and a toy dataset.
Figure 3: Overview of our approach. Our prompt includes repeated data example sets consisting of feature names and class-balanced groups, with the feature name at the end serving as a trigger for the LLM to generate realistic synthetic tabular data. The proposed unique variable remaps categorical values to distinct alphanumeric strings, ensuring clear distinction and variability among variables.
2 Method
In this study, we explore various prompt design options to maximize the in-context learning capability of LLMs for generating high-quality tabular data. Our goal is to propose an optimal approach that reliably and efficiently produces realistic tabular data to enhance ML classification performance, especially in the context of class imbalance. We start with a tabular dataset $T\in\mathds{R}^{n\times m}$, comprising $n$ instances, each with $m$ variables. We aim to generate a synthetic dataset, $\hat{T}\in\mathds{R}^{n^{\prime}\times m}$, which accurately mimics the characteristics of $T$, including class conditions, feature correlations, and value ranges. To achieve this, we explore several key prompt design elements: Data format, class presentation methods, variable mapping, and task specification.
We evaluate these design choices based on the quality of the generated data, with a primary focus on ML classification performance and generation efficiency, as detailed in Section 3. Our carefully crafted prompt, as depicted in Fig.3, begins with optional variable descriptions followed by multiple sets of data samples. Each set consists of feature names and samples grouped by target class. By adding feature names at the end of the prompt as a trigger, the LLM generates realistic data that improves ML performance and maintains feature correlations similar to the original data. Prompt examples and an illustration of the ablated versions are provided in Appendices I.5 and I.1, respectively.
2.1 Data format: Sentence vs CSV style
We investigate effective data formatting methods to represent tabular data within prompts, ensuring that LLMs can interpret the demonstrations and generate new data accurately. Tabular data can be formatted as plain text or as comma-separated values (CSV style). Let the feature name of the $k$-th variable be $v_{k}$, and the $i$-th observation of this variable be $o_{k,i}$. According to previous work [4], this value can be transformed into a sentence-style representation as $[v_{k},``is",o_{k,i},``,"]$. In contrast, the CSV style can be expressed as $[o_{k,i},``,"]$, presenting only values, with variable names for all variables specified as $\bm{v}=[v_{1},``,",v_{2},``,",\cdots,v_{m}]$ at the beginning of the data sample presentation. Here, the sentence style redundantly uses variable names and “is" tokens for every value, leading to higher token usage and computational cost than the CSV style. The CSV-style format allows for a higher volume of in-context learning examples within the same token constraints. Thus, our method utilizes a CSV-style format to maximize the number of examples, as providing a large number of data samples is crucial for ensuring sufficient representation of the original dataset in in-context learning.
2.2 Class presentation
Class presentation methods aim to ensure that all target classes are adequately represented within a prompt, particularly addressing class imbalance for minority classes.
Single-class vs. Multi-class generation    When generating data with class conditions, a primary consideration is whether to generate data for one class at a time or for all classes simultaneously. Generating data for a single class allows the model to focus on specific class characteristics, whereas generating data for multiple classes simultaneously enables comparison and contrast across classes. Our experiments indicate that generating data for multiple classes simultaneously, in a structured manner (explained subsequently), produces samples that more accurately represent each class compared to single-class generation. This suggests that generating data with an awareness of the characteristics of other classes is more effective.
Original vs. Balanced class ratio    When constructing a prompt to generate multiple classes simultaneously, naive random sampling from the original tabular dataset $T$ often results in an imbalanced ratio of samples for each target class. To address this, we consider balancing the sample numbers for minor and major classes rather than adhering to the original data distribution. Our findings indicate that balancing the sample numbers significantly improves the quality of generated data for minor classes, demonstrating that training with balanced data helps LLMs learn the true distributions of target classes more effectively.
Listing vs. Class grouping    Within a prompt, the data examples can either be listed in the order they are sampled without class distinction or grouped by class. Empirical evidence suggests that grouping by class leads to a better representation of class-specific feature correlations in the generated data. This grouping emphasizes contrasts between different classes and similarities within each group. As a result, the LLM generates more distinct samples for each class and more coherent samples within each class. Additionally, group-wise prompting can effectively condition LLMs to generate data with specific values. Without grouping, it is challenging to constrain LLMs to generate samples with specific attributes, often requiring extensive iterations until the desired samples are produced, which is inefficient. However, when examples are provided in groups by class, LLMs learn this and generate samples for each class in groups. Thus, grouping is crucial for conditioning the LLM to generate data with specific values and conditions. Moreover, grouping can be extended to condition on any arbitrary features other than target classes by grouping samples based on specific requirements.
2.3 Variable mapping: Original values vs. Unique variable mapping
As with other LLM-based approaches [4], our approach allows minimal data preprocessing, maintaining the original feature names and values. This significantly reduces the need for complex data manipulations often required in existing methods [18, 34] and helps to preserve the integrity of the raw data. However, one critical consideration arises when a dataset includes a large number of categorical variables that share the same values (e.g., extensive boolean variables). In this case, the data examples in the prompt can become monotonous and repetitive, making it challenging for the model to distinguish between different variables. This can significantly degrade the quality of the generated data. Moreover, when the LLM receives such input, the number of data samples generated at once can significantly decrease, and the generation efficiency can greatly decline.
To address this, we propose a unique variable mapping method, which remaps the values of each categorical variable to distinct random alphanumeric strings, as depicted in Fig. 3. Uniform substitution is applied across the dataset, maintaining the integrity of the data structure while introducing necessary variation. For example, consider a categorical variable $v_{1}$ with values of ‘t’ and ‘f’. Our approach remaps these values to distinct three-character strings, such as ‘JY0’ and ‘GGN,’ respectively, and applies them consistently across all data points in the dataset. Other categorical variables are similarly remapped to unique values. The transformed dataset is then used as in-context learning demonstrations. Although these new values do not hold inherent semantic meaning, they effectively represent categorical distinctions as symbols, allowing the LLM to identify and utilize patterns within them. This approach ensures that each variable has a unique representation, making the variables clearly distinguishable and contributing to more accurate and efficient data generation.
2.4 Task specification
A crucial aspect of the prompt design is to clearly define the task within the prompt, guiding the LLM to generate synthetic data as intended. To achieve this, the following options can be considered:
Explicit instruction vs. Completion triggering    We can clarify the task to the LLM by including explicit instructions to “generate new data samples." However, crafting specific instructions can be challenging, as LLMs are sensitive to minor variations in the prompt, making this approach complex for the tabular data generation task. A simpler and more effective method is providing input patterns that LLM can mimic to generate new data, only using data samples without explicit instruction. For example, let the $i$-th set of data samples for all $c$ classes be formatted as $\bm{S}_{i}=[\bm{v},\bm{s}_{i,1},\cdots,\bm{s}_{i,c}]$, where $\bm{s}_{i,j}$ consists of $n$ data samples for the $j$-th class. Here, the feature name $\bm{v}$ works as a header to indicate the start of a set. We input $t$ such sets in a prompt, $[\bm{S}_{1},\bm{S}_{2},\ldots,\bm{S}_{t}]$, creating a pattern and expecting the LLM to generate $\bm{S}_{t\texttt{+}1}=[\bm{v},\bm{s}_{t\texttt{+}1,1},\cdots,\bm{s}_{t\texttt{+}1,c}]$. To ensure this, we add a trigger to induce the LLM to complete the sequence. That is, we include the header of the new set $[\bm{v}]$ at the end of the prompt, which cues the LLM to generate the rest, $[\bm{s}_{t\texttt{+}1,1},\cdots,\bm{s}_{t\texttt{+}1,c}]$, learning the pattern from the input data that these groups are followed by $\bm{v}$. Thus, by creating a pattern in the input prompt, we can condition the LLM to generate a specific number ($c\times n$) of samples for each target class.
Providing contextual information    Optionally, including contextual information in the prompt can help the LLM understand the data structure and generate accurate samples. Therefore, if available, we add line-by-line variable descriptions at the beginning of the prompt. Additionally, one can consider clarifying the target class values at the beginning of each group. However, we observe that treating the target variable as one of the attributes and placing it as the first variable yields better performance, helping the LLM generate more coherent and accurate samples.
2.5 Overview of tabular data generation process
Synthetic data generation starts with creating a prompt template (see Appendix I.5 for examples). This template determines the target class and the number of sets, groups, and samples per group to construct demonstrations. If numerous categorical variables share the same unique values, we apply unique variable mapping to the dataset. Next, we randomly sample data examples from the dataset to fill in this template. Based on the provided examples, the LLM generates new data, which is then converted back into a tabular format. Instances with categorical variable values not present in the training data are discarded to maintain data authenticity. This process is repeated until the desired number of samples is obtained. Each iteration involves newly sampled examples from the original dataset, allowing the LLM to encounter various combinations and produce diverse samples. Although each individual generation is based on the distribution of the given subset, cumulatively, the iterations enable the model to explore the entire range of the original data. This method ensures that the output generated by the LLM comprehensively reflects the original data characteristics (See Appendix G).
Table 1: Comparison of machine learning classification performance when synthetic data are added to the original data. The results are averaged across four classifiers, with each model run five times. Complete results, including all baselines and standard deviation values, are provided in Appendix K.
Dataset Method #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
Travel Original - 58.12 (0.00) 71.00 (0.00) 57.00 (0.00) 85.00 (0.00)
+TVAE [30] +1K 59.78 (+1.66) 72.35 (+1.35) 62.00 (+5.00) 82.69 (-2.31)
+CTAB-GAN+ [34] +1K 54.66 (-3.46) 68.62 (-2.38) 53.00 (-4.00) 84.23 (-0.77)
+GReaT [4] +1K 60.95 (+2.83) 72.86 (+1.86) 58.80 (+1.80) 86.92 (+1.92)
+TabDDPM [18] +1K 53.20 (-4.92) 67.70 (-3.30) 50.40 (-6.60) 85.00 (0.00)
+Ours +1K  66.65 (+8.53) 78.23 (+7.23) 78.00 (+21.00) 78.46 (-6.54)
Sick Original - 87.81 (0.00) 91.22 (0.00) 82.83 (0.00) 99.61 (0.00)
+TVAE [30] +1K 87.77 (-0.04) 91.47 (+0.25) 83.37 (+0.54) 99.56 (-0.05)
+CTAB-GAN+ [34] +1K 82.35 (-5.46) 86.28 (-4.94) 72.83 (-10.00) 99.74 (+0.13)
+GReaT [4] +1K 87.23 (-0.58) 90.83 (-0.39) 82.07 (-0.76) 99.60 (-0.01)
+TabDDPM [18] +1K 85.17 (-2.64) 89.30 (-1.92) 79.02 (-3.81) 99.57 (-0.04)
+Ours +1K  88.71 (+0.90) 92.93 (+1.71) 86.41 (+3.58) 99.44 (-0.17)
HELOC Original - 71.01 (0.00) 73.21 (0.00) 67.89 (0.00) 78.52 (0.00)
+TVAE [30] +1K 71.12 (+0.11) 73.25 (+0.04) 68.15 (+0.26) 78.34 (-0.18)
+CTAB-GAN+ [34] +1K 71.03 (+0.02) 73.15 (-0.06) 68.13 (+0.24) 78.17 (-0.35)
+GReaT [4] +1K 70.35 (-0.66) 72.96 (-0.25) 66.22 (-1.67) 79.70 (+1.18)
+TabDDPM [18] +1K 70.65 (-0.36) 72.89 (-0.32) 67.51 (-0.38) 78.26 (-0.26)
+Ours +1K  71.92 (+0.91) 73.66 (+0.45) 69.96 (+2.07) 77.35 (-1.17)
Income Original - 66.90 (0.00) 76.45 (0.00) 57.28 (0.00) 95.61 (0.00)
+TVAE [30] +20K 66.96 (+0.06) 76.80 (+0.35) 59.13 (+1.85) 94.48 (-1.13)
+CTAB-GAN+ [34] +20K 66.49 (-0.41) 76.42 (-0.03) 58.14 (+0.86) 94.70 (-0.91)
+GReaT [4] +20K 67.95 (+1.05) 77.51 (+1.06) 60.69 (+3.41) 94.33 (-1.28)
+TabDDPM [18] +20K 66.85 (-0.05) 76.50 (+0.05) 57.70 (+0.42) 95.30 (-0.31)
+Ours +20K  69.16 (+2.26) 79.15 (+2.70) 66.45 (+9.17) 91.85 (-3.76)
Diabetes Original - 54.87 (0.00) 42.07 (0.00) 60.00 (0.00) 60.73 (0.00)
+TVAE [30] +10K 54.79 (-0.08) 41.96 (-0.11) 59.96 (-0.04) 60.71 (-0.02)
+CTAB-GAN+ [34] +10K 54.24 (-0.63) 41.52 (-0.55) 59.63 (-0.37) 60.01 (-0.72)
+GReaT [4] +10K 54.78 (-0.09) 41.98 (-0.09) 59.98 (-0.02) 60.61 (-0.12)
+TabDDPM [18] +10K 54.64 (-0.23) 41.83 (-0.24) 59.91 (-0.09) 60.55 (-0.18)
+Ours +10K  54.94 (+0.07) 42.14 (+0.07) 60.04 (+0.04) 60.82 (+0.09)
Thyroid Original - 94.23 (0.00) 95.08 (0.00) 91.14 (0.00) 99.02 (0.00)
+TVAE [30] +1K 90.45 (-3.78) 92.20 (-2.88) 86.36 (-4.78) 98.04 (-0.98)
+CTAB-GAN+ [34] +1K 27.46 (-66.77) 58.07 (-37.01) 16.14 (-75.00) 100.0 (+0.98)
+GReaT [4] +1K 91.31 (-2.92) 92.46 (-2.62) 85.91 (-5.23) 99.02 (0.00)
+TabDDPM [18] +1K 94.39 (+0.16) 96.26 (+1.18) 95.45 (+4.31) 97.06 (-1.96)
+Ours +1K  94.80 (+0.57) 96.39 (+1.31) 95.23 (+4.09) 97.55 (-1.47)
3 Experiments
Datasets    We use six real-world classification datasets from diverse domains to evaluate our method: Travel, Sick, HELOC, Income, Diabetes, and Thyroid. The Thyroid dataset is released after the training cut-off date for GPT-3.5-turbo-0613, which provides a more rigorous validation of our approach on the completely unseen dataset. For binary classification datasets, the minor class is set to be one. All duplicate data samples are removed from the datasets. Subsequently, we randomly split them into 80% training and 20% test sets. More details for the datasets are provided in Appendix I.2.
Evaluation measure    We evaluate our generation methodology from three critical aspects: ML classification performance, data generation success rate and stability with LLM in-context learning, and validity of generated data in terms of feature correlation and value range. For the classification task, we measure the F1 score, sensitivity, specificity, and balanced accuracy (BAL ACC). To measure the feature correlation, we use Pearson correlation and Cramér’s V correlation.
Baseline models    In our study, we compare our method with various generative models for tabular data generation, including TVAE [30], CopulaGAN [21], CTGAN [30], CTAB-GAN [33], CTAB-GAN+ [34], GReaT [4], and TabDDPM [18]. This comprehensive evaluation assesses the effectiveness of our approach across different data synthesis paradigms. To compare the prompting method, we also use the prompts used in CuratedLLM [25] and LITO [31].
Experimental details    Our method utilizes the OpenAI GPT-3.5 models, the GPT-3.5-turbo-0613 and GPT-3.5-turbo-16k-0613 versions, Mistral-7b-v0.1 [14], and Llama-2-7b [27]. Unless stated otherwise, four machine learning classification models are used: XGBoost [8], CatBoost [23], LightGBM [16], and Gradient boosting classifier [11]. We execute five independent runs for each model and report the average values. Additional details are provided in Appendix I.
Table 2: Comparison of classification performance using synthetic data generated by open-source LLMs, Llama2 and Mistral, with the proposed prompting strategy, utilizing the Gradient Boosting classifier. We report the average performances for five runs. #syn denotes the number of synthetic samples integrated into the original dataset.
Dataset Method Model #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
Travel Original - - 60.00$\pm$0.00 72.31$\pm$0.00 60.00$\pm$0.00 84.62$\pm$0.00
+Ours Mistral +1K 66.67$\pm$0.00 78.00$\pm$0.00 76.00$\pm$0.00 80.00$\pm$0.00
+Ours Llama2 +1K 67.80$\pm$0.00 79.23$\pm$0.00 80.00$\pm$0.00 78.46$\pm$0.00
+Ours GPT3.5 +1K 70.18$\pm$0.00 80.77$\pm$0.00 80.00$\pm$0.00 81.54$\pm$0.00
Sick Original - - 84.09$\pm$0.00 89.86$\pm$0.00 80.43$\pm$0.00 99.28$\pm$0.00
+Ours Mistral +1K 88.42$\pm$0.00 95.15$\pm$0.00 91.30$\pm$0.00 99.00$\pm$0.00
+Ours Llama2 +1K 85.53$\pm$0.95 93.14$\pm$0.52 87.39$\pm$0.97 98.88$\pm$0.06
+Ours GPT3.5 +1K 91.95$\pm$0.00 93.41$\pm$0.00 86.96$\pm$0.00 99.86$\pm$0.00
HELOC Original - - 71.32$\pm$0.04 73.39$\pm$0.03 68.47$\pm$0.06 78.31$\pm$0.00
+Ours Mistral +1K 71.48$\pm$0.00 73.74$\pm$0.00 68.00$\pm$0.00 79.47$\pm$0.00
+Ours Llama2 +1K 70.77$\pm$0.00 73.08$\pm$0.00 67.37$\pm$0.00 78.79$\pm$0.00
+Ours GPT3.5 +1K 71.93$\pm$0.02 73.68$\pm$0.02 69.90$\pm$0.00 77.45$\pm$0.04
3.1 Machine learning classification performance using the synthetic data
We assess the quality of the synthetic data samples by comparing the machine learning classification performance averaged over 20 experiments. As shown in Table 1, our method achieves state-of-the-art F1 scores and balanced accuracy across all six datasets, surpassing the baselines that require model training. Notably, our method is the only one that consistently outperforms the original data in both F1 score and balanced accuracy across all six datasets. Other methods often result in worse performance than the original data, indicating that the data generated by baseline methods disrupts the original data distribution. These findings underscore the robustness and effectiveness of our method, affirming its superiority over baseline models, even for challenging datasets.
Moreover, baseline methods such as GReaT and TabDDPM exhibit significantly lower sensitivity compared to specificity, particularly for Travel and Income. Their high balanced accuracy is due to high specificity, but their severely low sensitivity indicates a failure to generate appropriate minority class data. In contrast, our method significantly improves sensitivity and achieves a balance between sensitivity and specificity while also attaining the best-balanced accuracy and F1 score compared to the baselines. For example, in Travel, our method achieves a sensitivity of 78%, which is 19%p higher than the second-best method, GReaT, at 58.8%. In Income, our method shows 66.45% sensitivity, surpassing the second-best score of 60.69%. These results highlight the effectiveness of our method in accurately representing minority classes in tabular datasets.
Overall, our findings demonstrate that our approach is generally applicable across various imbalanced tabular datasets and excels in generating high-quality samples that improve ML classification performance. For detailed confusion matrix results, see Appendix J, and for further analyses exploring the impact of adding synthetic data only to the minor class or using synthetic data alone, see Appendix A.
3.2 Open-source LLMs
We also apply our prompting method to open-source LLMs, including Llama2 [27] and Mistral [14]. As shown in Table 2, our method exhibits robust performance when used with open-source LLMs across various datasets. Notably, in the Sick dataset, Mistral demonstrates the highest balanced accuracy, indicating its effectiveness in generating high-quality tabular data. These results validate the broader applicability and general effectiveness of our approach with open-source LLMs.
Table 3: Comparison of F1 scores with ablated methods using the Gradient Boosting classifier. Asterisk (*) in Income indicates where only $\texttt{+}1$K synthetic samples are used due to a low success rate in sample generation. The ‘-’ symbol denotes where our unique variable mapping is not needed.
Format Class Balance Group Unique Sick ($\texttt{+}1$K) Travel ($\texttt{+}1$K) Income ($\texttt{+}20$K)
Sentence single 81.52$\pm$0.65 43.24$\pm$0.00 69.62$\pm$0.00
CSV-style single 86.60$\pm$0.00 58.67$\pm$0.00 70.49$\pm$0.00
Sentence multi 77.95$\pm$0.33 58.33$\pm$0.00 67.96$\pm$0.00
CSV-style multi 87.74$\pm$0.96 50.98$\pm$0.00 68.56$\pm$0.00
Sentence multi 81.68$\pm$0.49 55.56$\pm$0.00 69.26$\pm$0.00
CSV-style multi 84.68$\pm$0.38 70.37$\pm$0.00 70.64$\pm$0.00
CSV-style multi 86.60$\pm$0.00 - -
CSV-style multi 81.55$\pm$0.00 70.18$\pm$0.00 70.17$\pm$0.00
CSV-style multi 91.95$\pm$0.00 - -
Table 4: Comparison of token efficiency and generation stability on the Income dataset.
Format Class Balance Group Unique #set Input tokens $\downarrow$ Output samples $\uparrow$ Success rate $\uparrow$
Sentence single - 1832.9 1.26 96%
CSV-style single - 1024.6 1.11 16%
Sentence multi - 1827.4 0.97 97%
CSV-style multi - 912.8 0.38 38%
Sentence multi - 1938.7 1.73 97%
CSV-style multi - 920.3 0.88 31%
CSV-style multi 1 965.6 8.6 48%
CSV-style multi 2 1014.6 10.52 94%
Table 5: Comparison of classification performance on task specification elements in prompt design, averaged across four classifiers: XGBoost, CatBoost, LightGBM, and Gradient Boosting classifier. Methods marked with an asterisk (*) use prompts as proposed in the respective papers.
Method #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
Instruction-$\text{CuratedLLM}^{*}$ +1K 15.86$\pm$5.13 56.02$\pm$3.54 17.17$\pm$11.80 94.86$\pm$4.93
Instruction-$\text{LITO}^{*}$ +1K 21.32$\pm$1.65 72.06$\pm$2.49 84.46$\pm$3.10 59.66$\pm$3.50
Ours w/ class distinction +1K 74.06$\pm$2.64 96.23$\pm$0.98 96.74$\pm$1.93 95.72$\pm$0.60
Ours w/o var description +1K 76.06$\pm$4.56 96.29$\pm$1.40 96.41$\pm$2.76 96.18$\pm$1.06
Ours +1K 88.71$\pm$1.98 92.93$\pm$0.91 86.41$\pm$1.85 99.44$\pm$0.27
Figure 4: The difference between Cramér’s V correlation matrices of real and synthetic datasets for categorical variables in the Travel dataset. More intense colors indicate larger differences, with positive differences shown in red and negative differences shown in blue. Black indicates where the correlation is not measured since only one unique value is generated for those variables.
3.3 Exploring optimal prompt design through ablation studies
We investigate key prompt design elements focusing on ML performance and generation efficiency, as detailed in Tables 3 and Table 4. Token efficiency is measured by the number of input tokens required for a fixed number of input samples. LLM generation efficiency is evaluated based on the average number of samples generated per iteration and the success rate of generating at least one valid data sample per inference. This evaluation is based on 100 iterations, each with 20 random input samples.
Our results indicate that CSV-style prompting generally outperforms sentence-style prompting in ML performance with the same number of input samples. While generating data for one class at a time yields good results for the Sick and Income datasets, it results in a low F1 score of 58.67% on the Travel dataset. However, generating data for multiple classes simultaneously with balanced samples improves this score to 70.37%. Furthermore, applying unique variable mapping consistently enhances F1 scores on the Sick dataset. The combined approach of random mapping and grouping achieves higher F1 score of 91.95% compared to 86.60% with random mapping but without grouping, underscoring the positive impact of this combined approach. These findings indicate that prompt design significantly influences the quality of generated data, emphasizing the need for carefully tailored prompts to generate high-quality synthetic tabular data.
For the Travel and Income datasets, performance without grouping is slightly better than with grouping, but the differences are within 0.5%p, making their performance on par. However, these show significant differences in LLM generation efficiency. Without grouping, an average of 0.88 samples are generated per iteration, whereas grouping increases this to an average of 8.6 samples per iteration. Repeating the set twice within the prompt further improves generation efficiency, boosting the average output to over 10 samples and the success rate to 94%, by creating input patterns that LLMs can mimic to generate data. Unique variable mapping also significantly enhances LLM generation efficiency and success rates, as detailed in Appendix B.
Furthermore, we conduct an ablation study on the task specification elements, as shown in Table 5. Our approach of triggering completion significantly enhances overall ML performance compared to the instruction-based prompting methods. Interestingly, adding class distinctions within the prompt achieves the best sensitivity, indicating that the distinction facilitates better generation for minority classes. Providing variable descriptions improves balanced accuracy but reduces the F1 score, suggesting that this element may be optional rather than a definitive improvement. For a comparison of generation efficiency using these methods, see Appendix B.
3.4 Analysis on feature correlation
We compare the feature correlations of the categorical variables between original and synthetic data, as illustrated in Fig. 4. Our method, utilizing Mistral and Llama2, produces data that most closely mirrors the original feature correlations for both minor and major classes. Moreover, when compared to the ablated versions, our grouping method significantly contributes to preserving feature correlation for the minority class. For additional feature correlation results on the Sick dataset, see Appendix C.
4 Conclusion
This study demonstrates the effectiveness of using LLMs for tabular data generation and introduces an optimal solution for prompt design to generate realistic and reliable data without additional training, particularly addressing class imbalance. Our extensive evaluations demonstrate the robustness and effectiveness of our method, which significantly improves ML classification performance on the six real-world tabular datasets. This highlights its practicality and potential impact on improving model performance in various application scenarios, contributing significantly to the field.
Appendix
This supplementary material enhances the main manuscript by providing detailed experimental results and additional visualizations.
  • Appendix A presents more experimental results on the machine learning classification for imbalanced classes.
  • Appendix B offers complete additional ablation study results.
  • Appendix C provides feature correlation results for the Sick dataset.
  • Appendix D studies the impact of sample size on the machine learning performance.
  • Appendix E comprehensively analyzes the experiments conducted on the toy dataset.
  • Appendix F discusses about the broader impacts of our research.
  • Appendix G outlines limitations of our work.
  • Section H provides related work.
  • Appendix I provides comprehensive information on the datasets, comparison baselines, implementation details, and prompt design and examples.
  • Appendix J presents confusion matrix results, highlighting the effectiveness of our method in addressing minor class problems.
  • Appendix K shows the complete results of Table 1.
Appendix A Additional experimental results
Our experiments with synthetic tabular data for imbalanced classes follow three main approaches: (1) Adding synthetic data to both minor and major classes of the existing dataset and evaluating machine learning classification performance. The results of this approach are presented in Table 1 of the main manuscript. (2) Adding synthetic data only to the minor class, with detailed results provided in Appendix A.1. (3) Evaluating classification performance using only synthetic data without the original dataset, as discussed in Appendix A.2.
Table 6: Comparison of classification performance for imbalanced data by generating synthetic data for the minority class to match the size of the majority class. The average performances of the Gradient Boosting Classifier over five runs are reported.
Dataset Method #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
Travel Original - 60.00$\pm$0.00 72.31$\pm$0.00 60.00$\pm$0.00 84.62$\pm$0.00
+SMOTE +163 63.45$\pm$1.05 75.08$\pm$0.69 68.00$\pm$0.00 82.15$\pm$1.38
+SMOTENC +163 62.61$\pm$2.54 74.74$\pm$2.00 70.40$\pm$3.58 79.08$\pm$0.84
+TVAE +163 54.55$\pm$0.00 68.46$\pm$0.00 60.00$\pm$0.00 76.92$\pm$0.00
+CopulaGAN +163 52.46$\pm$0.00 66.62$\pm$0.00 64.00$\pm$0.00 69.23$\pm$0.00
+CTGAN +163 57.63$\pm$0.00 70.92$\pm$0.00 68.00$\pm$0.00 73.85$\pm$0.00
+GReaT +163 62.07$\pm$0.00 74.46$\pm$0.00 72.00$\pm$0.00 76.92$\pm$0.00
+Ours +163 72.13$\pm$0.00 83.23$\pm$0.00 88.00$\pm$0.00 78.46$\pm$0.00
Income Original - 69.05$\pm$0.00 78.02$\pm$0.00 61.03$\pm$0.00 95.00$\pm$0.00
+SMOTE +13,487 69.62$\pm$0.25 79.84$\pm$0.23 68.86$\pm$0.57 90.81$\pm$0.14
+SMOTENC +13,487 69.79$\pm$0.11 80.51$\pm$0.06 71.80$\pm$0.23 89.22$\pm$0.19
+TVAE +13,487 68.26$\pm$0.00 79.12$\pm$0.00 68.43$\pm$0.00 89.82$\pm$0.00
+CopulaGAN +13,487 67.76$\pm$0.02 81.12$\pm$0.01 80.15$\pm$0.03 82.09$\pm$0.00
+CTGAN +13,487 68.51$\pm$0.00 78.82$\pm$0.00 66.26$\pm$0.00 91.38$\pm$0.00
+GReaT +13,487 70.04$\pm$0.00 83.36$\pm$0.00 85.08$\pm$0.00 81.64$\pm$0.00
+Ours +13,487 71.09$\pm$0.02 83.60$\pm$0.01 83.53$\pm$0.03 83.66$\pm$0.00
A.1 Comparative analysis with adding the new dataset only to the minor class
One of the major advantages of our method is its ability to generate high-quality synthetic data for minor classes, even with limited samples. Similar to oversampling techniques like SMOTE, we add synthetic dataset only to the minor class and evaluate machine learning classification performance. To balance the ratio of major to minor classes, we add only minor class samples to the original data, then measure the classification performance of the machine learning model. For this task, we additionally investigate the effectiveness of our method compared to SMOTE and SMOTENC [6]. As shown in Table 6, adding minor class samples generated from our method leads to higher F1 scores and balanced accuracy on both Travel and Income datasets, compared to when samples generated by other baselines are added. In the Travel dataset, where the number of minor class samples is substantially low, baseline models, especially, CopulaGAN and CTGAN, demonstrate markedly low sensitivity, indicating their failure to accurately learn the minor class data distribution. In contrast, our method demonstrates its efficacy in producing high-quality data for minor classes on the Travel dataset. Compared to the Original, it significantly improves the F1 score by 12.13%p, underscoring its superior performance of our approach.
A.2 Comparative analysis of classification performance using only synthetic data
Table 7: Comparison of classification performance using only synthetic data for training. We report average values for five runs employing Gradient Boosting Classifier.
Dataset Method #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
Travel Original - 60.00$\pm$0.00 72.31$\pm$0.00 60.00$\pm$0.00 84.62$\pm$0.00
TVAE 1K 46.51$\pm$0.00 63.85$\pm$0.00 40.00$\pm$0.00 87.69$\pm$0.00
CopulaGAN 1K 0.00$\pm$0.00 49.23$\pm$0.00 0.00$\pm$0.00 98.46$\pm$0.00
CTGAN 1K 12.50$\pm$0.00 50.15$\pm$0.00 8.00$\pm$0.00 92.31$\pm$0.00
GReaT 1K 62.50$\pm$0.00 73.85$\pm$0.00 60.00$\pm$0.00 87.69$\pm$0.00
Ours 1K 70.00$\pm$0.00 81.23$\pm$0.00 84.00$\pm$0.00 78.46$\pm$0.00
HELOC Original - 71.32$\pm$0.04 73.39$\pm$0.03 68.47$\pm$0.06 78.31$\pm$0.00
TVAE 1K 68.91$\pm$0.10 71.50$\pm$0.08 65.28$\pm$0.12 77.72$\pm$0.07
CopulaGAN 1K 69.50$\pm$0.00 70.82$\pm$0.00 69.17$\pm$0.00 72.47$\pm$0.00
CTGAN 1K 68.71$\pm$0.02 71.98$\pm$0.03 63.36$\pm$0.00 80.60$\pm$0.05
GReaT 1K 64.14$\pm$0.10 69.54$\pm$0.07 55.76$\pm$0.15 83.33$\pm$0.11
Ours 1K 71.48$\pm$0.06 70.17$\pm$0.05 78.90$\pm$0.16 61.44$\pm$0.16
We compare the classification performance using only synthetic data for training, as shown in Table 7. For both the Travel and HELOC datasets, the classification model trained on synthetic datasets generated by our model outperform the one trained on original data in terms of F1 score and Sensitivity. This improvement in performance suggests that the data generated by our method accurately represents the distribution of the original data. Furthermore, it highlights the effectiveness of our class-balanced generation approach in improving the performance of classification models.
Table 8: Comparison of token efficiency and generation stability for ablating prompt design elements on the Sick dataset
Format Class Balance Group Unique #set Input tokens $\downarrow$ Output samples $\uparrow$ Success rate $\uparrow$
Sentence single - 3867.2 0.52 52%
CSV-style single - 1223.6 2.42 38%
Sentence multi - 3879.9 0.52 52%
CSV-style multi - 1226.3 3.05 48%
Sentence multi - 3870.0 0.63 63%
CSV-style multi - 1117.3 0.70 16%
CSV-style multi - 1985.4 3.00 15%
CSV-style multi 1 1331.3 4.73 64%
CSV-style multi 2 1439.9 6.93 99%
CSV-style multi 2 2060.8 17.68 95%
Table 9: Comparison of token efficiency and generation stability for task specification in prompt design on the Sick dataset.
Method Input tokens $\downarrow$ Output samples $\uparrow$ Success rate $\uparrow$
Instruction-$\text{CuratedLLM}^{*}$ 4677.4 21.67 95%
Instruction-$\text{LITO}^{*}$ 1277.4 4.06 57%
Ours w/ class distinction $2$.7 9.7 91%
Ours w/o var description 2066.9 15.96 88%
Ours 2060.8 17.68 95%
Appendix B Additional ablation study results
This section presents additional ablation study results on the Sick datasets. We evaluate the token efficiency and generation stability of ablated versions based on 100 iterations, each with 20 random input samples. As shown in Table 8, for the Sick dataset, due to its repetitive monotonous values, the number of generated samples is low, even with grouping. Using a unique variable mapper in such cases significantly increases the number of samples generated per attempt from 6.93 to 17.68. This improvement demonstrates that our approach effectively controls the quantity of generated data samples, ensuring the stable production of synthetic tabular data. Moreover, the ability to generate multiple data samples in a single inference significantly can lowers generation time and costs.
We also compare the efficiency and stability of generation when ablating task specification elements in our prompt design, as shown in Table 9. The prompt design from CuratedLLM [25], despite requiring the highest number of input tokens, achieves high performance in terms of output samples and success rate. However, the generated data shows relatively low quality, resulting in a low classification performance with an F1 score and balanced accuracy, as discussed in Section 3.3. In contrast, the prompt design of LITO [31] is the most efficient in terms of input tokens but produces the fewest output samples and has the lowest success rate, indicating instability in synthetic data generation.
Our proposed method balances efficiency and performance, using approximately half the input tokens compared to CuratedLLM’s prompt while generating a high number of output samples and maintaining a high success rate. Additionally, the synthetic data generated by our method achieves the highest classification performance in terms of F1 score, as shown in Section 3.3. When class distinctions are provided, the average number of output samples decreased from 17.68 to 9.70. This result indicates that it is more robust to distinguish groups with simple numbering, such as A and B, rather than explicitly specifying class information in each group. Including variable descriptions slightly reduces the number of output samples and success rate compared to not including them, but the difference is not significant. This suggests that the grouping strategy plays a more crucial role in robust data generation than the presence of variable descriptions.
Figure 5: The differences between Pearson correlation matrices of real data and synthetic data for numerical variables in the Sick dataset. More intense colors indicate larger differences, with positive differences shown in red and negative differences shown in blue.
Appendix C Analysis of feature correlation between numerical variables
In the main manuscript, we visualize the feature correlation between categorical variables. Here, we analyze the Pearson correlation between numerical variables in the Sick dataset and compare the differences in feature correlation values to the original data, as shown in Fig. 5. As a result, our methodology consistently shows feature correlations similar to the original data for both minor and major classes. In contrast, other baselines either show significant deviations in feature correlation for all classes or perform well for major classes but not for minor ones. These results demonstrate that our methodology shows superior performance compared to the baselines in generating high-quality data that accurately represents both minor and major classes within the original dataset.
Appendix D Impacts of sample size on the classification performance
Figure 6: Impact of sample size on the classification performance on the Income and Diabetes datasets when classifiers are trained exclusively on synthetic samples generated from our proposed method. Black dashed lines denote the F1 score using only the original data.
Figure 7: Complete generation results on an imbalanced toy dataset. Major class is shown in blue and minor class in red. Our approach via in-context learning of LLM achieves (1) clear class boundaries, (2) accurate feature correlations, (3) proper value ranges, (4) well-represented relationships between categorical (last row) and numerical variables, and (5) better diversity compared to the fine-tuned GReaT [4] model.
We investigate how varying the number of synthetic samples affects classification performance. This analysis is conducted on the Income and Diabetes datasets, and the results are shown in Fig. 6. For the Income dataset, sample sizes range from 500 to 10,000, while for Diabetes, the range was 500 to 20,000. For both datasets, a consistent increase in the F1 score is observed as the number of samples grow. This result demonstrates that our method can adequately represent the original data distribution by generating a sufficient number of synthetic data samples.
Figure 8: The illustration of the input samples (blue and orange) and the generated samples (green and red) for each inference on the imbalanced toy dataset when the our proposed approach received 180 input samples. Empty spaces in the visualization indicate instances where the model failed to generate samples. To produce a total of 1000 samples, the model performed 23 inferences.
Figure 9: The illustration of the input samples (blue and orange) and the generated samples (green and red) for each inference on the imbalanced toy dataset when the ablated model without grouping received 180 input samples. Empty spaces in the visualization indicate instances where the model failed to generate samples. To produce a total of 1,000 samples, the model performed 37 inferences.
Appendix E Complete qualitative results on toy dataset
In Fig. 7, we illustrate the complete qualitative results on the toy dataset. We assume a challenging multivariate data scenario in an imbalanced binary classification situation, where variables with and without correlations are mixed. Using the scikit-learn library  [22], we created a total of 12 variables. These variables have strong correlations in pairs. There are 11 numerical variables and 1 categorical variable, and the categorical variable is shown on the x-axis in the last row of Fig. 7. In the figure, the major class is shown in blue, and the minor class is shown in red.
The generated results can be evaluated from four aspects. Firstly, whether each data sample is correctly generated to belong to each class, i.e., whether the class boundary is clear. Secondly, whether the shapes formed by the correlation between the variables, such as circles or moons, are well represented. Thirdly, whether the values generated fall within the acceptable range for each variable. Finally, whether the relationship between the categorical and numerical variables is well represented.
The baseline method, GReaT [4], has successfully learned from the training data to generate class-separated samples. However, upon closer inspection, there are samples belonging to different classes. Ours, despite generating samples without training, shows similar or better quality.
Overall, CSV-style generation produces better class separation boundaries compared to sentence-style generation. Regarding class presentation, giving a single class can sometimes result in good generation outcomes, but it fails to accurately capture the correlation with the major class, resulting in shapes like squares instead of circles, and it also fails to accurately generate the categorical variable. On the other hand, when generating multi-class samples, if the sample numbers are not balanced, the generation results for the minor class can be compromised. Even when balanced, providing grouping leads to clearer class distinctions. Additionally, adding unique variable mapping significantly improves the performance, resulting in a well-distinguished categorical variable.
In addition, we visualize the inputs and outputs of both our method and the ablated model without grouping, as shown in Fig. 8 and Fig. 9. For each inference, 180 class-balanced samples are input into the LLM prompt. To generate a total of 1,000 samples, our method required 23 inferences, whereas the ablated model needed 37 inferences, demonstrating the stability of our method in generating synthetic data samples. In terms of data generation quality, our method successfully generates both major and minor classes in most inferences, with the distribution of the generated samples closely mirroring that of the actual input data. In contrast, the ablated model exhibits a higher number of inferences where the generated data deviates from the input data distribution.
Appendix F Broader Impact
Our method is generally applicable to any tabular classification task with minimal pre-processing, making it accessible for various research and industrial applications. The potential positive impacts include enhanced data accessibility and privacy preservation, which can improve decision-making processes in fields like healthcare and finance. Additionally, it reduces the need for using real data, thereby protecting individual privacy. However, there are potential negative impacts, such as the risk of data misuse. Synthetic data generation could be misused for malicious purposes, such as creating fake datasets to deceive systems or individuals.
Figure 10: The illustration of the input (blue and orange) and generated (green and red) samples from our method where input samples are exclusively taken from either the left or right side of the circle.
Appendix G Limitations
If the input samples do not adequately represent the data, the LLM may fail to generate samples that follow the original data distribution, leading to degraded performance. As shown in Fig. 10, if the input samples do not fully capture the original data shape, the generated data may fail to accurately follow this shape. A limitation of our work is that our current method relies on providing a large number of data samples to ensure sufficient representation in in-context learning. Future work will focus on increasing efficiency by selectively choosing key samples that can help the LLM generate data effectively with fewer examples.
Appendix H Related work
In the field of tabular data synthesis, traditional methods have predominantly focused on sampling techniques, such as SMOTE [6], which have been foundational in addressing class imbalances in tabular data. Although useful, these methods often fall short in capturing complex inter-feature correlations [2] and typically lack the versatility needed for regression tasks or addressing multiple feature imbalances. Recently, deep-learning-based generative models have demonstrated their efficacy in generating realistic tabular data [7, 9, 15, 30, 33, 34]. For example, CTAB-GAN [33] introduced a generator that mitigates the problem caused by imbalanced discrete variables. Additionally, diffusion models have also exhibited promising results in tabular data generation. TabDDPM [18] introduced a diffusion process for both numerical and categorical features in tabular data for accurate data generation. Large language model (LLM)-based methods have emerged as a recent trend in tabular dataset generation. GReaT [4] surpassed previous methods by fine-tuning LLMs to generate tabular data formatted in natural text. However, this fine-tuning process is resource-intensive, requiring extensive training of large models for each dataset.
As an alternative, leveraging the in-context learning capabilities of LLMs presents a promising solution. LLMs have demonstrated the ability to solve various problems, learn and mimic patterns from given input data, and generate new data without additional training for parameter updates [5, 13, 17, 19, 28, 29, 32]. Despite this potential, there is limited research on what LLMs generate for synthetic tabular data and what the optimal prompts for high-quality tabular data generation would be. Previous studies [31, 25] have shown the potential of LLMs for in-context learning of tabular data generation. However, these studies have not sufficiently explored optimal prompts for synthetic tabular data generation, leading to instability in the generated samples. This instability necessitates additional validation modules and a two-stage structure.
Given that the performance of in-context learning with LLMs heavily depends on prompt design [10, 26, 28, 32], carefully designing prompts specifically for imbalanced tabular data is essential to maximize LLM capabilities in effectively generating realistic data. Recognizing the importance of optimal prompt design and the potential instability in generated data, we propose a novel method for generating high-quality synthetic tabular data through tailored prompting. The proposed method is generally applicable to various types of tabular datasets without the need for training, thereby offering a cost-effective solution in data synthesis.
Appendix I Additional experimental details
This section provides the illustration of ablated versions (Appendix I.1), comprehensive information on the datasets (Appendix I.2) and comparison baselines (Appendix I.3), implementation details (Appendix I.4), and final prompt design and examples (Appendix I.5).
I.1 Ablation prompt examples
This section visualizes five ablated versions of our proposed prompt design, as shown in Fig. 11. The ablations involve removing or altering four key components: data format, balanced class ratio, multi-class generation, and class grouping.
Figure 11: Illustration of ablated versions.
I.2 Dataset
The detailed description of the datasets are shown in Table 10. The variable descriptions are directly sourced from dataset platforms like Kaggle or the UCI repository. In our approach, we designate the class with fewer instances as the positive class for binary classification tasks.
Table 10: Details of the datasets used to evaluate synthetic tabular data generation models.
Dataset #Class #categorical features #numerical features #samples Domain
Travel 2 4 2 447 Marketing
Sick 2 21 6 3,711 Medical
HELOC 2 - 23 9,872 Financial
Income 2 8 6 32,561 Social
Diabetes 3 36 11 101,766 Medical
Thyroid 2 16 1 364 Medical
I.3 Baseline reproducibility
In our research, we utilize implementations from the Synthetic Data Vault [21] for TVAE [30], CTGAN [30], and CopulaGAN, while for the GReaT [4] model, we employ code from its official repository. In our experiment, the GReaT model fails to accurately generating column names ‘max_glu_serum’ and ‘A1Cresult’ in the Diabetes dataset, even after training for 85 epochs that reported in the original paper. Extending the training to 105 epochs does not rectify this issue, with ‘max_glu_serum’ and ‘A1Cresult’ being correctly generated in only approximately 2.8% and 12.1% of instances, respectively. For these columns, failure cases are treated as NaN value.
We utilize the official open-source code from the official TabDDPM GitHub repository to reproduce CTAB-GAN [33], CTAB-GAN+ [34], and TabDDPM [18]. For the Adult Income and California datasets, the hyperparameters provided are used. For other datasets lacking provided hyperparameters, we apply the hyperparameters from the Wilt dataset, which, according to the corresponding paper, demonstrated the largest performance improvement over the original data.
Due to the lack of official code provided by the authors of LITO [31], we adopt the prompt design described in their original paper. Specifically, we use the prompt for LITO-C to generate minor class data. To generate samples for both majority and minority classes, we also use the modified prompt to generate majority class data. Similarly, for CuratedLLM [25], we follow the prompt design detailed in their paper. Although LITO and CuratedLLM propose multi-stage techniques for data generation, we focus solely on evaluating their prompt designs.
I.4 Experimental setup
For classification, we utilize the target class of each dataset to divide groups. Specifically, the Diabetes dataset has three groups, and the Travel, Sick, HELOC, Income and Thyroid datasets have two groups. For the process of unique variable mapping, each discrete value is transformed into a combination of three characters, including uppercase letters and digits.
Regarding hyperparameter settings of downstream machine learning models: For the Gradient Boosting Classifier, we use the default hyperparameters provided by scikit-learn library. For XGBoost [8], CatBoost [23], and LightGBM [16], we conduct 5-fold cross-validation on the training set of the original data, optimizing two key hyperparameters: learning rate and max depth. These optimized settings are then consistently applied across all experiments, aligning the model’s performance closely with the characteristics of the original data. Where possible, we train the classification models using a single NVIDIA GeForce RTX 3090 GPU on an Ubuntu 18.04 system.
I.5 More details on the prompt design
This section provides an in-depth explanation of the construction of the proposed prompt. Prompt examples for generating synthetic data are provided in Tables 11-16.
Altogether, the final prompt template for synthetic data generation consists of four parts.
Descriptions (Optional). The prompt may start with descriptions of the variables in the dataset, provided line-by-line, if precise and available. These descriptions define what each column in the dataset represents, offering contextual understanding to the LLM about the data. It will be located at the top of the prompt (but not repeated to maintain token efficiency) for additional context.
Set-level examples. Then, a set-level examples are present. This section begins with a data entry format header, specifying the order and names of each feature. This header acts as a guide for the LLM, indicating the start of an example set, which is crucial for interpreting the structure of the tabular data. Following the header, there’s a line specifying a group name, signaling the start of a specific group. Instances belonging to this group are then listed, one per line, in a CSV style. This process is repeated for each group within the set.
Repetition of set-Level template. To reinforce the data structure and patterns, this set-level template is repeated several times within the prompt, each time with distinct data samples. This repetition is key in enabling the LLM to recognize, learn, and subsequently replicate the set format in its synthetic data generation process.
Trigger for LLM data generation. After all the example sets are presented, the same header that indicates the start of an example set is included at the end of the prompt. This header serves as a signal for the LLM to commence generating synthetic data, maintaining the structure and patterns established in the prompt.
By following this structured approach, the LLM is guided to generate synthetic data that not only adheres to the format and structure of the original dataset but also incorporates variability and distinctiveness in the data patterns. The inclusion of descriptions, careful structuring of set-level examples, and strategic repetition of templates all contribute to the effectiveness of this synthetic data generation method.
Appendix J Confusion matrix results
This section provides confusion matrix results on four binary classification datasets with Gradient Boosting classifier. The confusion matrix in Fig. 12 reveals that machine learning classification model classifies most accurately when the data generated from our method is added to the original data.
Figure 12: Confusion matrix on the Travel, Sick, HELOC, and Income datasets using Gradient Boosting classifier.
Table 11: Example of a data synthesis prompt for the Travel dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data.
Template Prompt sample Descriptions Churn: whether customer churns or doesnt churn for tour and travels company,
age: the age of customer,
FrequentFlyer: whether customer takes frequent flights,
AnnualIncomeClass: class of annual income of user,
ServicesOpted: number of times services opted during recent years,
AccountSyncedToSocialMedia: whether company account of user synchronised to their social media,
[…]
Set Header Churn,Age,FrequentFlyer,AnnualIncomeClass,ServicesOpted, AccountSyncedToSocialMedia,BookedHotelOrNot Group A.
Churn,28,Yes,High Income,6,No,Yes
Churn,37,Yes,Low Income,4,Yes,Yes
[…]
Group B.
Doesnt churn,38,No,Low Income,1,Yes,No
Doesnt churn,28,No Record,Low Income,5,No,Yes
[…]
Set Header Churn,Age,FrequentFlyer,AnnualIncomeClass,ServicesOpted, AccountSyncedToSocialMedia,BookedHotelOrNot Group A.
Churn,30,Yes,High Income,4,No,No
Churn,28,No,Low Income,6,No,Yes
[…]
Group B.
Doesnt churn,37,Yes,Low Income,1,No,No
Doesnt churn,36,No,Middle Income,1,No,Yes
[…]
Set Header Churn,Age,FrequentFlyer,AnnualIncomeClass,ServicesOpted, AccountSyncedToSocialMedia,BookedHotelOrNot Group A.
Churn,27,Yes,High Income,5,No,No
Churn,37,No,Low Income,5,Yes,No
[…]
Group B.
Doesnt churn,37,Yes,Low Income,1,No,No
Doesnt churn,36,No,Middle Income,1,No,Yes
[…]
Set Header (trigger) Churn,Age,FrequentFlyer,AnnualIncomeClass,ServicesOpted, AccountSyncedToSocialMedia,BookedHotelOrNot
Table 12: Example of a data synthesis prompt for the Sick dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data. In this scenario, our unique variable mapping is employed, whereby each unique value of a variable is consistently substituted with a unique three-character alphanumeric string. This approach ensures diversity and robustness in the synthesized data.
Template Prompt sample Descriptions Class: hypothyroidism is a condition in which the thyroid gland is underperforming or producing too little thyroid hormone,
age: the age of an patient,
sex: the biological sex of an patient,
TSH: thyroid stimulating hormone,
T3: triiodothyronine hormone,
TT4: total levothyroxine hormone,
T4U: levothyroxine hormone uptake,
FTI: free levothyroxine hormone index,
[…]
Set Header Class,age,sex,on_thyroxine,query_on_thyroxine, on_antithyroid_medication,sick,pregnant,thyroid_surgery, I131_treatment,query_hypothyroid,query_hyperthyroid,lithium, goitre,tumor,hypopituitary,psych,TSH_measured,TSH,T3_measured,T3, TT4_measured,TT4,T4U_measured,T4U,FTI_measured,FTI,referral_source Group A.
[…]
Group B.
[…]
Set Header Class,age,sex,on_thyroxine,query_on_thyroxine, on_antithyroid_medication,sick,pregnant,thyroid_surgery, I131_treatment,query_hypothyroid,query_hyperthyroid,lithium, goitre,tumor,hypopituitary,psych,TSH_measured,TSH,T3_measured,T3, TT4_measured,TT4,T4U_measured,T4U,FTI_measured,FTI,referral_source Group A.
[…]
Group B.
[…]
Set Header (trigger) Class,age,sex,on_thyroxine,query_on_thyroxine, on_antithyroid_medication,sick,pregnant,thyroid_surgery, I131_treatment,query_hypothyroid,query_hyperthyroid,lithium, goitre,tumor,hypopituitary,psych,TSH_measured,TSH,T3_measured,T3, TT4_measured,TT4,T4U_measured,T4U,FTI_measured,FTI,referral_source
Table 13: Example of a data synthesis prompt for the HELOC dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data.
Template Prompt sample Descriptions Set Header RiskPerformance,ExternalRiskEstimate,MSinceOldestTradeOpen, MSinceMostRecentTradeOpen,AverageMInFile,NumSatisfactoryTrades, NumTrades60Ever2DerogPubRec,NumTrades90Ever2DerogPubRec, PercentTradesNeverDelq,MSinceMostRecentDelq, MaxDelq2PublicRecLast12M,MaxDelqEver,NumTotalTrades, NumTradesOpeninLast12M,PercentInstallTrades, MSinceMostRecentInqexcl7days,NumInqLast6M,NumInqLast6Mexcl7days, NetFractionRevolvingBurden,NetFractionInstallBurden, NumRevolvingTradesWBalance,NumInstallTradesWBalance, NumBank2NatlTradesWHighUtilization,PercentTradesWBalance Group A.
Bad,90,211,6,102,17,0,0,100,-7,7,8,17,1,0,0,1,1,0,-8,0,-8,0,0
[…]
Group B.
[…]
Set Header RiskPerformance,ExternalRiskEstimate,MSinceOldestTradeOpen, MSinceMostRecentTradeOpen,AverageMInFile,NumSatisfactoryTrades, NumTrades60Ever2DerogPubRec,NumTrades90Ever2DerogPubRec, PercentTradesNeverDelq,MSinceMostRecentDelq, MaxDelq2PublicRecLast12M,MaxDelqEver,NumTotalTrades, NumTradesOpeninLast12M,PercentInstallTrades, MSinceMostRecentInqexcl7days,NumInqLast6M,NumInqLast6Mexcl7days, NetFractionRevolvingBurden,NetFractionInstallBurden, NumRevolvingTradesWBalance,NumInstallTradesWBalance, NumBank2NatlTradesWHighUtilization,PercentTradesWBalance Group A.
Bad,63,150,3,50,29,3,3,91,75,6,3,32,4,75,-7,6,6,33,90,1,7,1,73
[…]
Group B.
[…]
Set Header (trigger) RiskPerformance,ExternalRiskEstimate,MSinceOldestTradeOpen, MSinceMostRecentTradeOpen,AverageMInFile,NumSatisfactoryTrades, NumTrades60Ever2DerogPubRec,NumTrades90Ever2DerogPubRec, PercentTradesNeverDelq,MSinceMostRecentDelq, MaxDelq2PublicRecLast12M,MaxDelqEver,NumTotalTrades, NumTradesOpeninLast12M,PercentInstallTrades, MSinceMostRecentInqexcl7days,NumInqLast6M,NumInqLast6Mexcl7days, NetFractionRevolvingBurden,NetFractionInstallBurden, NumRevolvingTradesWBalance,NumInstallTradesWBalance, NumBank2NatlTradesWHighUtilization,PercentTradesWBalance
Table 14: Example of a data synthesis prompt for the Income dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data.
Template Prompt sample Descriptions Set Header income,age,workclass,fnlwgt,education,education_num, marital-status,occupation,relationship,race,sex,capital-gain, capital-loss,hours-per-week,native-country Group A.
>50K,52,Private,298215,Bachelors,13,Married-civ-spouse, Craft-repair,Husband,White,Male,0,0,50,United-States
>50K,30,Self-emp-inc,321990,Masters,14,Married-civ-spouse, Exec-managerial,Husband,White,Male,15024,0,60,?
[…]
Group B.
<=50K,42,Private,572751,Preschool,1,Married-civ-spouse, Craft-repair,Husband,White,Male,0,0,40,Nicaragua
<=50K,25,Self-emp-not-inc,159909,Assoc-voc,11,Married-civ-spouse, Farming-fishing,Husband,White,Male,0,0,40,United-States
[…]
Set Header income,age,workclass,fnlwgt,education,education_num, marital-status,occupation,relationship,race,sex,capital-gain, capital-loss,hours-per-week,native-country Group A.
>50K,38,Private,58108,Bachelors,13,Married-civ-spouse, Exec-managerial,Husband,White,Male,0,0,50,United-States
>50K,31,State-gov,124020,Assoc-acdm,12,Married-civ-spouse, Tech-support,Husband,White,Male,0,0,40,United-States
[…]
Group B.
<=50K,51,Private,138514,Assoc-voc,11,Divorced,Tech-support, Unmarried,Black,Female,0,0,48,United-States
<=50K,18,Private,205218,11th,7,Never-married,Sales,Own-child,White, Female,0,0,20,United-States
[…]
Set Header (trigger) income,age,workclass,fnlwgt,education,education_num, marital-status,occupation,relationship,race,sex,capital-gain, capital-loss,hours-per-week,native-country
Table 15: Example of a data synthesis prompt for the Diabetes dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data. In this scenario, our unique variable mapping is employed, whereby each unique value of a variable is consistently substituted with a unique three-character alphanumeric string. This approach ensures diversity and robustness in the synthesized data.
Template Prompt sample Descriptions Set Header readmitted,encounter_id,patient_nbr,race,gender,age,weight, admission_type_id,discharge_disposition_id,admission_source_id, time_in_hospital,payer_code,medical_specialty,num_lab_procedures, num_procedures,num_medications,number_outpatient,number_emergency, number_inpatient,diag_1,diag_2,diag_3,number_diagnoses, max_glu_serum,A1Cresult,metformin,repaglinide,nateglinide, chlorpropamide,glimepiride,acetohexamide,glipizide,glyburide, tolbutamide,pioglitazone,rosiglitazone,acarbose,miglitol, troglitazone,tolazamide,insulin,glyburide-metformin, glipizide-metformin,glimepiride-pioglitazone, metformin-rosiglitazone,metformin-pioglitazone,change,diabetesMed Group A.
AD6,39850434,428733,HX2,LL7,10F,H7H,2,7,1,13,QTD,3S2,44,4,15,1,0, 6,QDY,HR7,8K9,9,JAT,N0H,K06,2TV,65H,A7C,MK7,JWO,HXK,SVQ,BMY,FCV, ZRU,IDJ,8CO,78A,NXL,CWD,6TQ,VDM,HMH,0OU,HNM,NGY,NC4
[…]
Group B.
YRB,163732050,91571517,7MT,LL7,BM0,H7H,3,1,1,1,QTD,SKI,17,6,9,0,0, 1,6RA,4UR,0TR,9,JAT,N0H,K06,2TV,65H,A7C,MK7,JWO,HXK,SVQ,BMY,FCV, ZRU,IDJ,8CO,78A,NXL,PDO,6TQ,VDM,HMH,0OU,HNM,UXM,R51
[…]
Group C.
PS6,274193592,68737500,7MT,LL7,10F,H7H,1,3,7,5,QBR,SKI,60,1,18,0,0, 0,DL8,AL8,P2T,9,JAT,ZAX,K06,2TV,65H,A7C,MK7,JWO,HXK,SVQ,BMY,FCV, ZRU,IDJ,8CO,78A,NXL,CWD,6TQ,VDM,HMH,0OU,HNM,NGY,NC4
[…]
Table 16: Example of a data synthesis prompt for the Thyroid dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data.
Template Prompt sample Descriptions Set Header Recurred,Age,Gender,Smoking,Hx Smoking,Hx Radiothreapy, Thyroid Function,Physical Examination,Adenopathy,Pathology, Focality,Risk,T,N,M,Stage,Response Group A.
Yes,46,M,Yes,No,No,Euthyroid,Single nodular goiter-left, Bilateral,Follicular,Uni-Focal,High,T4b,N1b,M1,II,Structural Incomplete
Yes,27,M,No,No,No,Euthyroid,Multinodular goiter,Bilateral, Papillary,Multi-Focal,Intermediate,T3a,N1b,M0,I, Structural Incomplete
[…]
Group B.
[…]
Set Header Recurred,Age,Gender,Smoking,Hx Smoking,Hx Radiothreapy, Thyroid Function,Physical Examination,Adenopathy,Pathology, Focality,Risk,T,N,M,Stage,Response Group A.
Yes,37,M,No,No,No,Euthyroid,Multinodular goiter,Bilateral, Papillary,Multi-Focal,Intermediate,T3a,N1b,M0,I,Structural Incomplete
Yes,63,M,Yes,No,No,Euthyroid,Single nodular goiter-right, Right,Papillary,Multi-Focal,Intermediate,T3a,N1b,M0,II,Structural Incomplete
[…]
Group B.
[…]
Set Header (trigger) Recurred,Age,Gender,Smoking,Hx Smoking,Hx Radiothreapy, Thyroid Function,Physical Examination,Adenopathy,Pathology, Focality,Risk,T,N,M,Stage,Response
Appendix K Complete results on comparison with baseline methods
Table K offers the complete performance comparison results with standard deviation values of 20 independent experiments on all six datasets.
Table 17: Complete results on comparison of classification performance improvement using synthetic data from baseline models and our method, averaged across four classifiers: XGBoost, CatBoost, LightGBM, and Gradient Boosting classifier, with each model run five times. “#syn” represents the number of synthetic samples added to the original data.
Dataset Method #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
Travel Original - 58.12$\pm$2.04 (0.00) 71.00$\pm$1.41 (0.00) 57.00$\pm$3.40 (0.00) 85.00$\pm$0.68 (0.00)
+TVAE [30] +1K 59.78$\pm$4.89 (+1.66) 72.35$\pm$3.52 (+1.35) 62.00$\pm$6.81 (+5.00) 82.69$\pm$1.31 (-2.31)
+CopulaGAN [21] +1K 21.76$\pm$2.00 (-36.36) 55.52$\pm$0.83 (-15.48) 12.80$\pm$1.64 (-44.20) 98.23$\pm$1.89 (+13.23)
+CTGAN [30] +1K 29.84$\pm$3.61 (-28.28) 57.79$\pm$1.56 (-13.21) 19.20$\pm$2.46 (-37.80) 96.38$\pm$1.15 (+11.38)
+CTAB-GAN [33] +1K 56.07$\pm$8.29 (-2.05) 69.58$\pm$5.14 (-1.42) 51.00$\pm$7.88 (-6.00) 88.15$\pm$3.31 (+3.15)
+CTAB-GAN+ [34] +1K 54.66$\pm$3.64 (-3.46) 68.62$\pm$2.40 (-2.38) 53.00$\pm$3.40 (-4.00) 84.23$\pm$2.05 (-0.77)
+GReaT [4] +1K 60.95$\pm$2.59 (+2.83) 72.86$\pm$1.80 (+1.86) 58.80$\pm$3.69 (+1.80) 86.92$\pm$0.79 (+1.92)
+TabDDPM [18] +1K 53.20$\pm$4.10 (-4.92) 67.70$\pm$2.69 (-3.30) 50.40$\pm$4.19 (-6.60) 85.00$\pm$1.31 (0.00)
 +Ours +1K  66.65$\pm$2.53 (+8.53) 78.23$\pm$2.10 (+7.23) 78.00$\pm$4.59 (+21.00) 78.46$\pm$2.50 (-6.54)
Sick Original - 87.81$\pm$2.46 (0.00) 91.22$\pm$0.95 (0.00) 82.83$\pm$1.71 (0.00) 99.61$\pm$0.22 (0.00)
+TVAE [30] +1K 87.77$\pm$2.88 (-0.04) 91.47$\pm$1.33 (+0.25) 83.37$\pm$2.47 (+0.54) 99.56$\pm$0.23 (-0.05)
+CopulaGAN [21] +1K 83.60$\pm$1.33 (-4.21) 86.61$\pm$0.53 (-4.61) 73.37$\pm$0.97 (-9.46) 99.86$\pm$0.10 (+0.25)
+CTGAN [30] +1K 87.52$\pm$1.68 (-0.29) 89.86$\pm$0.52 (-1.36) 79.89$\pm$0.97 (-2.94) 99.82$\pm$0.19 (+0.21)
+CTAB-GAN [33] +1K 86.12$\pm$1.49 (-1.69) 89.51$\pm$1.24 (-1.71) 79.35$\pm$2.49 (-3.48) 99.68$\pm$0.06 (+0.07)
+CTAB-GAN+ [34] +1K 82.35$\pm$4.16 (-5.46) 86.28$\pm$2.49 (-4.94) 72.83$\pm$4.86 (-10.00) 99.74$\pm$0.14 (+0.13)
+GReaT [4] +1K 87.23$\pm$1.87 (-0.58) 90.83$\pm$1.09 (-0.39) 82.07$\pm$2.10 (-0.76) 99.60$\pm$0.12 (-0.01)
+TabDDPM [18] +1K 85.17$\pm$2.07 (-2.64) 89.30$\pm$1.18 (-1.92) 79.02$\pm$2.26 (-3.81) 99.57$\pm$0.10 (-0.04)
 +Ours +1K  88.71$\pm$1.98 (+0.90) 92.93$\pm$0.91 (+1.71) 86.41$\pm$1.85 (+3.58) 99.44$\pm$0.27 (-0.17)
HELOC Original - 71.01$\pm$0.47 (0.00) 73.21$\pm$0.31 (0.00) 67.89$\pm$0.82 (0.00) 78.52$\pm$0.34 (0.00)
+TVAE [30] +1K 71.12$\pm$0.32 (+0.11) 73.25$\pm$0.33 (+0.04) 68.15$\pm$0.21 (+0.26) 78.34$\pm$0.48 (-0.18)
+CopulaGAN [21] +1K 71.23$\pm$0.27 (+0.22) 73.32$\pm$0.25 (+0.11) 68.37$\pm$0.31 (+0.48) 78.26$\pm$0.36 (-0.26)
+CTGAN [30] +1K 70.82$\pm$0.24 (-0.19) 73.06$\pm$0.27 (-0.15) 67.60$\pm$0.25 (-0.29) 78.52$\pm$0.55 (0.00)
+CTAB-GAN [33] +1K 70.60$\pm$0.41 (-0.41) 72.87$\pm$0.21 (-0.34) 67.39$\pm$0.97 (-0.50) 78.36$\pm$0.74 (-0.16)
+CTAB-GAN+ [34] +1K 71.03$\pm$0.05 (+0.02) 73.15$\pm$0.10 (-0.06) 68.13$\pm$0.21 (+0.24) 78.17$\pm$0.38 (-0.35)
+GReaT [4] +1K 70.35$\pm$0.33 (-0.66) 72.96$\pm$0.24 (-0.25) 66.22$\pm$0.49 (-1.67) 79.70$\pm$0.21 (+1.18)
+TabDDPM [18] +1K 70.65$\pm$0.18 (-0.36) 72.89$\pm$0.14 (-0.32) 67.51$\pm$0.35 (-0.38) 78.26$\pm$0.30 (-0.26)
 +Ours +1K  71.92$\pm$0.11 (+0.91) 73.66$\pm$0.17 (+0.45) 69.96$\pm$0.21 (+2.07) 77.35$\pm$0.51 (-1.17)
Income Original - 66.90$\pm$2.12 (0.00) 76.45$\pm$1.48 (0.00) 57.28$\pm$3.41 (0.00) 95.61$\pm$0.46 (0.00)
+TVAE [30] +20K 66.96$\pm$1.36 (+0.06) 76.80$\pm$1.11 (+0.35) 59.13$\pm$2.92 (+1.85) 94.48$\pm$0.71 (-1.13)
+CopulaGAN [21] +20K 66.75$\pm$1.72 (-0.15) 76.73$\pm$1.33 (+0.28) 59.16$\pm$3.41 (+1.88) 94.29$\pm$0.81 (-1.32)
+CTGAN [30] +20K 66.22$\pm$0.82 (-0.68) 76.27$\pm$0.65 (-0.18) 57.95$\pm$1.73 (+0.67) 94.59$\pm$0.44 (-1.02)
+CTAB-GAN [33] +20K 66.48$\pm$2.11 (-0.42) 76.31$\pm$1.51 (-0.14) 57.45$\pm$3.61 (+0.17) 95.17$\pm$0.58 (-0.44)
+CTAB-GAN+ [34] +20K 66.49$\pm$1.14 (-0.41) 76.42$\pm$0.90 (-0.03) 58.14$\pm$2.41 (+0.86) 94.70$\pm$0.63 (-0.91)
+GReaT [4] +20K 67.95$\pm$1.36 (+1.05) 77.51$\pm$1.05 (+1.06) 60.69$\pm$2.56 (+3.41) 94.33$\pm$0.51 (-1.28)
+TabDDPM [18] +20K 66.85$\pm$1.83 (-0.05) 76.50$\pm$1.34 (+0.05) 57.70$\pm$3.26 (+0.42) 95.30$\pm$0.60 (-0.31)
 +Ours +20K  69.16$\pm$1.01 (+2.26) 79.15$\pm$0.82 (+2.70) 66.45$\pm$1.98 (+9.17) 91.85$\pm$0.49 (-3.76)
Diabetes Original - 54.87$\pm$1.37 (0.00) 42.07$\pm$1.23 (0.00) 60.00$\pm$0.64 (0.00) 60.73$\pm$1.63 (0.00)
+TVAE [30] +10K 54.79$\pm$1.40 (-0.08) 41.96$\pm$1.24 (-0.11) 59.96$\pm$0.66 (-0.04) 60.71$\pm$1.65 (-0.02)
+CopulaGAN [21] +10K 54.27$\pm$1.48 (-0.60) 41.59$\pm$1.26 (-0.48) 59.73$\pm$0.75 (-0.27) 59.97$\pm$1.65 (-0.76)
+CTGAN [30] +10K 54.72$\pm$1.13 (-0.15) 41.92$\pm$1.02 (-0.15) 59.86$\pm$0.53 (-0.14) 60.63$\pm$1.32 (-0.10)
+CTAB-GAN [33] +10K 54.22$\pm$1.28 (-0.65) 41.53$\pm$1.09 (-0.54) 59.73$\pm$0.59 (-0.27) 59.91$\pm$1.46 (-0.82)
+CTAB-GAN+ [34] +10K 54.24$\pm$1.16 (-0.63) 41.52$\pm$1.01 (-0.55) 59.63$\pm$0.57 (-0.37) 60.01$\pm$1.31 (-0.72)
+GReaT [4] +10K 54.78$\pm$1.31 (-0.09) 41.98$\pm$1.17 (-0.09) 59.98$\pm$0.60 (-0.02) 60.61$\pm$1.55 (-0.12)
+TabDDPM [18] +10K 54.64$\pm$1.50 (-0.23) 41.83$\pm$1.28 (-0.24) 59.91$\pm$0.75 (-0.09) 60.55$\pm$1.78 (-0.18)
 +Ours +10K  54.94$\pm$1.43 (+0.07) 42.14$\pm$1.29 (+0.07) 60.04$\pm$0.66 (+0.04) 60.82$\pm$1.68 (+0.09)
Thyroid Original - 94.23$\pm$1.99 (0.00) 95.08$\pm$1.60 (0.00) 91.14$\pm$3.12 (0.00) 99.02$\pm$1.01 (0.00)
+TVAE [30] +1K 90.45$\pm$1.89 (-3.78) 92.20$\pm$1.65 (-2.88) 86.36$\pm$3.30 (-4.78) 98.04$\pm$0.00 (-0.98)
+CopulaGAN [21] +1K 86.73$\pm$3.99 (-7.50) 88.71$\pm$2.91 (-6.37) 78.41$\pm$5.08 (-12.73) 99.02$\pm$1.01 (0.00)
+CTGAN [30] +1K 76.40$\pm$6.58 (-17.83) 81.14$\pm$4.55 (-13.94) 62.27$\pm$9.10 (-28.87) 100.0$\pm$0.00 (+0.98)
+CTAB-GAN [33] +1K 53.49$\pm$3.71 (-40.74) 68.30$\pm$1.73 (-26.78) 36.59$\pm$3.45 (-54.55) 100.0$\pm$0.00 (+0.98)
+CTAB-GAN+ [34] +1K 27.46$\pm$7.80 (-66.77) 58.07$\pm$2.60 (-37.01) 16.14$\pm$5.21 (-75.00) 100.0$\pm$0.00 (+0.98)
+GReaT [4] +1K 91.31$\pm$1.61 (-2.92) 92.46$\pm$0.99 (-2.62) 85.91$\pm$1.40 (-5.23) 99.02$\pm$1.01 (0.00)
+TabDDPM [18] +1K 94.39$\pm$1.09 (+0.16) 96.26$\pm$0.50 (+1.18) 95.45$\pm$0.00 (+4.31) 97.06$\pm$1.01 (-1.96)
 +Ours +1K  94.80$\pm$1.02 (+0.57) 96.39$\pm$0.62 (+1.31) 95.23$\pm$1.02 (+4.09) 97.55$\pm$0.87 (-1.47)

### v4

@@
-Exploring Prompting Methods for Mitigating Class Imbalance through Synthetic Data Generation with Large Language Models
+EPIC: Effective Prompting for Imbalanced-Class Data Synthesis in Tabular Data Classification via Large Language Models
 Abstract
-Large language models (LLMs) have demonstrated impressive in-context learning capabilities across various domains.
-Inspired by this, our study explores the effectiveness of LLMs in generating realistic tabular data to mitigate class imbalance.
-We investigate and identify key prompt design elements such as data format, class presentation, and variable mapping to optimize the generation performance.
-Our findings indicate that using CSV format, balancing classes, and employing unique variable mapping produces realistic and reliable data, significantly enhancing machine learning performance for minor classes in imbalanced datasets.
-Additionally, these approaches improve the stability and efficiency of LLM data generation.
-We validate our approach using six real-world datasets and a toy dataset, achieving state-of-the-art performance in classification tasks.
-The code is available here.
-Figure 1: Generation results on an imbalanced toy dataset.
-The major class is shown in blue, and the minor class in red.
-Our approach via in-context learning of LLM achieves (1) clear class boundaries, (2) accurate feature correlations, (3) proper value ranges, (4) well-represented relationships between categorical and numerical variables (last row), and (5) better diversity compared to the fine-tuned GReaT [4] model.
-Detailed explanations and complete results are provided in Appendix E.
+Large language models (LLMs) have demonstrated remarkable in-context learning capabilities across diverse applications.
+In this work, we explore the effectiveness of LLMs for generating realistic synthetic tabular data, identifying key prompt design elements to optimize performance.
+We introduce EPIC, a novel approach that leverages balanced, grouped data samples and consistent formatting with unique variable mapping to guide LLMs in generating accurate synthetic data across all classes, even for imbalanced datasets.
+Evaluations on real-world datasets show that EPIC achieves state-of-the-art machine learning classification performance, significantly improving generation efficiency.
+These findings highlight the effectiveness of EPIC for synthetic tabular data generation, particularly in addressing class imbalance.
+Our source code for our work is available here.
+Figure 1: Generation results on an imbalanced toy dataset with majority and minority classes.
+Our approach, leveraging in-context learning with LLMs, achieves (1) distinct class boundaries, (2) accurate feature correlations, (3) well-matched value ranges, (4) robust numerical-categorical relationships (last row), and (5) comprehensive data distribution coverage, with improvements over its ablated versions and the fine-tuned GReaT model [4].
+Complete results are available in Fig.
+9.
+Figure 2: Comparison of ML classification performance with synthetic data on the Travel dataset.
+Results are averaged across four classifiers: XGBoost, CatBoost, LightGBM, and gradient boosting classifier, with each classifier run five times.
+Our method uses the GPT-3.5-turbo model.
 1 Introduction
-Tabular data plays a crucial role in various classification tasks, such as disease diagnosis [1] and financial analysis [3, 24].
-However, challenges such as data scarcity and class imbalance, particularly for rare yet crucial events, can severely degrade the performance of machine learning (ML) models.
-To address these, synthetic data generation methods have been developed to augment tabular datasets cost-effectively.
-Traditional methods such as SMOTE and its variants [6, 12, 20] focus on generating data for minority classes to alleviate class imbalance.
-Recently, advanced deep learning-based generation models have shown promising results, such as TVAE [30], CTAB-GAN [33], and TabDDPM [18].
-Large language models (LLMs) have also demonstrated significant potential in generating realistic tabular data, showing robust performance across numerical and categorical data [4, 31].
-However, these models often require extensive, data-specific fine-tuning, which can lead to overfitting on major classes or dominant feature values when applied to small and imbalanced tabular datasets.
-An alternative approach leverages the in-context learning capabilities of LLMs.
-Extensive studies have demonstrated that LLMs can solve complex problems, learn and mimic patterns from input data, and generate new data without parameter updates [13, 17, 19, 28, 29, 32].
-This capability holds significant potential for generating realistic and reliable tabular data to address class imbalance.
-However, the in-context learning capabilities of LLMs largely depend on a prompt design [10, 26, 28, 32], while these methods are not directly applicable to tabular data generation, especially for imbalanced class scenarios.
-This is because tabular data is not inherently in plain text and involves unique challenges such as identifying feature correlations and accurately representing underrepresented attributes.
-Therefore, carefully crafted prompt methods specifically for imbalanced tabular data are essential to maximize LLM capabilities in generating realistic data effectively.
-While several studies have highlighted the potential of LLMs for generating synthetic tabular data using in-context learning [25, 31], they have not sufficiently explored various prompt design elements, such as data format or class presentation.
-This study aims to investigate the effectiveness of LLMs in resolving class imbalance problems and identify the optimal prompts to maximize this effectiveness.
-Figure 2: Comparison of ML classification performance when adding synthetic data to the Travel dataset.
-The results are averaged across four classifiers, XGBoost, CatBoost, LightGBM, and Gradient Boosting Classifier.
-Each classifier is run five times, resulting in a total of 20 experiments.
-To answer these questions, we explore various prompt methods and analyze the generated results to identify specific prompt structures and templates that significantly improve the quality of synthetic tabular data generated by LLMs, thereby enhancing ML classification performance.
-We propose several prompt design elements, including data format, class presentation methods, variable mapping, and task specification.
-Our findings indicate that providing data in CSV format, balancing and grouping classes, and applying a unique variable mapper to distinguish different variables significantly improve the quality and efficiency of synthetic data generation, underscoring their effectiveness in optimizing LLM performance.
+Tabular data, consisting of mixed variable types such as numerical and categorical variables, represents a widely applicable and essential data format [30].
+It plays a crucial role in enhancing decision-making and efficiency in various real-world applications, such as finance, healthcare, manufacturing, and natural sciences [1, 3, 26].
+However, challenges such as data scarcity and class imbalance, particularly for rare yet crucial events, often significantly degrade the performance of machine learning (ML) models, often resulting in reduced accuracy for underrepresented classes.
+To address these challenges, efficient synthetic data generation methods have been developed to augment tabular datasets.
+Traditional methods, such as synthetic minority oversampling technique (SMOTE) and its variants [5, 11, 21], focus on generating minority class samples to alleviate class imbalance.
+Recently, advanced generative models, such as TVAE [35], CTAB-GAN [39], and TabDDPM [19], have shown promising results in producing high-quality synthetic tabular data.
+Large language models (LLMs) have also demonstrated significant potential for generating realistic tabular data, effectively handling both numerical and categorical variables [4, 37].
+However, these models often require extensive, data-specific fine-tuning, which can increase the risk of overfitting to majority classes or dominant feature values, especially in small and highly imbalanced tabular datasets.
+An alternative approach involves leveraging the in-context learning capabilities of LLMs.
+Existing work has shown that LLMs can solve complex reasoning tasks, learn and mimic patterns from input data, and augment textual datasets without parameter updates [13, 20, 34].
+However, designing effective prompts to optimize this capability for tabular data generation is challenging, especially for imbalanced datasets.
+This is because tabular data is not naturally expressed in textual form and involves unique challenges, such as identifying feature correlations and accurately representing underrepresented attributes.
+Therefore, carefully crafted prompt design tailored to synthetic tabular data generation is essential to fully leverage the capabilities of LLMs.
+While several studies have employed in-context learning with LLMs for synthetic tabular data generation [27, 37], few have conducted comprehensive investigations into optimal prompt designs that significantly impact data quality and generation efficiency.
+In response, this study examines key components of prompt design and identifies an effective method for generating high-quality synthetic tabular data, particularly addressing class imbalance.
+We introduce a novel approach, EPIC, which leverages the in-context learning capabilities of LLMs to produce synthetic tabular data with balanced class representation.
+EPIC incorporates prompt design strategies such as CSV style formatting, balanced class grouping, and a unique variable mapper, which together contribute to generating synthetic data that accurately represents class-specific distributions and feature correlations, as illustrated in Fig.
+1.
+Extensive evaluations across six real-world datasets demonstrate that EPIC significantly improves the data quality and generation efficiency, achieving state-of-the-art performance.
 As shown in Fig.
-1, the data generated using this prompt design exhibit the most distinct boundaries between target classes, accurate feature correlation, and variable ranges.
-Moreover, this approach significantly improves the classification performance of ML models when applied to real-world datasets.
-As shown in Fig.
-2, baseline methods exhibit low sensitivity, indicating that they fail to accurately generate minority class samples due to inherent class biases.
-In contrast, our method overcomes these, achieving high sensitivity and balanced performance for all classes.
-In summary, our contributions are:
-  • Our study explores and analyzes the effectiveness of LLMs in generating realistic synthetic tabular data via in-context learning, offering guidelines for designing prompts tailored to effectively exploit LLMs for this task, particularly addressing class imbalance.
-  • We propose simple yet effective prompting methods that generate high-quality class-specific tabular data, maintaining feature correlation and value ranges.
-  • The proposed approach significantly improves ML classification performance and enhances data generation efficiency and stability via in-context learning.
-  • The proposed approach is generally applicable and easy to implement for any tabular classification task with minimal pre-processing, providing publicly available code.
-  • We validate the effectiveness of our approach through extensive evaluation across six real-world public tabular data classification datasets and a toy dataset.
+2, baselines exhibit low sensitivity, struggling to accurately generate minority class samples due to inherent class imbalances.
+In contrast, EPIC achieves high sensitivity and balanced performance for all classes, underscoring its robustness and practicality for real-world applications.
+In summary, our key contributions are as follows:
+  • Our study explores the effectiveness of LLMs in generating realistic synthetic tabular data through in-context learning, providing prompt design guidelines to efficiently generate high-quality data while addressing class imbalance.
+  • We propose EPIC, a simple yet effective prompting method that uses balanced, grouped data samples with unique variable mapping to generate tabular data that accurately represents both minority and majority classes, preserving feature correlations and overall data distribution.
+  • The proposed approach is model-agnostic, generally applicable to various LLMs, and easy to implement for any tabular data with minimal preprocessing requirements.
+  • Extensive experiments on six real-world public tabular datasets and one toy dataset demonstrate the effectiveness of our approach, significantly improving ML classification performance and data generation efficiency.
+By addressing class imbalance and enhancing classification outcomes, our work contributes to the advancement of the crucial field of tabular data research, significantly impacting various domains.
 Figure 3: Overview of our approach.
@@
 2 Method
-In this study, we explore various prompt design options to maximize the in-context learning capability of LLMs for generating high-quality tabular data.
-Our goal is to propose an optimal approach that reliably and efficiently produces realistic tabular data to enhance ML classification performance, especially in the context of class imbalance.
-We start with a tabular dataset $T\in\mathds{R}^{n\times m}$, comprising $n$ instances, each with $m$ variables.
-We aim to generate a synthetic dataset, $\hat{T}\in\mathds{R}^{n^{\prime}\times m}$, which accurately mimics the characteristics of $T$, including class conditions, feature correlations, and value ranges.
-To achieve this, we explore several key prompt design elements: Data format, class presentation methods, variable mapping, and task specification.
-We evaluate these design choices based on the quality of the generated data, with a primary focus on ML classification performance and generation efficiency, as detailed in Section 3.
-Our carefully crafted prompt, as depicted in Fig.3, begins with optional variable descriptions followed by multiple sets of data samples.
-Each set consists of feature names and samples grouped by target class.
-By adding feature names at the end of the prompt as a trigger, the LLM generates realistic data that improves ML performance and maintains feature correlations similar to the original data.
-Prompt examples and an illustration of the ablated versions are provided in Appendices I.5 and I.1, respectively.
+In this study, we investigate various prompt design components to maximize the in-context learning capabilities of LLMs for generating high-quality tabular data.
+Our objective is to develop an optimized approach that reliably and effectively produces realistic tabular data, accurately representing both minority and majority classes to improve ML classification performance, especially addressing class imbalance.
+Formally, given a tabular dataset $T$ of dimensions $n\times m$, with $n$ samples and $m$ variables, we aim to generate a synthetic dataset $\hat{T}$ of dimensions $n^{\prime}\times m$ that accurately reflects key characteristics of $T$, including class-specific attributes, feature correlations and, data distributions.
+To achieve this, we explore the following key prompt design elements: data format (Section 2.1), class presentation methods (Section 2.2), variable mapping (Section 2.3), and task specification (Section 2.4).
+These design choices are evaluated through extensive analyses on the public datasets from diverse domains, focusing on ML classification performance and generation efficiency.
+Detailed analysis is provided in Section 3, with Fig.
+12 in Appendix C.1 illustrating the ablated versions.
+Building on these analyses, we introduce EPIC, a structured prompting method designed to guide LLMs in synthetic data generation while effectively addressing class imbalance, as depicted in Fig.
+3.
+EPIC begins with optional variable descriptions, followed by a series of structured data sample sets, each containing balanced samples for each class.
+Each set consists of feature names and samples organized by target class.
+To prompt the LLM to generate a corresponding set of synthetic samples, feature names are appended at the end of the prompt to serve as a trigger, leveraging the pattern recognition capabilities of LLMs.
+The following sections detail each design component of EPIC.
 2.1 Data format: Sentence vs CSV style
@@
 Here, the sentence style redundantly uses variable names and “is" tokens for every value, leading to higher token usage and computational cost than the CSV style.
-The CSV-style format allows for a higher volume of in-context learning examples within the same token constraints.
+The CSV-style format enables a higher volume of in-context learning examples within the same token constraints.
 Thus, our method utilizes a CSV-style format to maximize the number of examples, as providing a large number of data samples is crucial for ensuring sufficient representation of the original dataset in in-context learning.
 2.2 Class presentation
-Class presentation methods aim to ensure that all target classes are adequately represented within a prompt, particularly addressing class imbalance for minority classes.
+Class presentation methods aim to ensure adequate representation of all target classes within a prompt, particularly addressing underrepresented classes in imbalanced datasets.
+The construction of data samples within the prompt is crucial, as it can either worsen or mitigate existing imbalances.
+To address this, we explore three prompt design options and propose a class balancing and grouping approach that enables accurate and representative generation of data samples across all target classes.
 Single-class vs.
-Multi-class generation    When generating data with class conditions, a primary consideration is whether to generate data for one class at a time or for all classes simultaneously.
-Generating data for a single class allows the model to focus on specific class characteristics, whereas generating data for multiple classes simultaneously enables comparison and contrast across classes.
-Our experiments indicate that generating data for multiple classes simultaneously, in a structured manner (explained subsequently), produces samples that more accurately represent each class compared to single-class generation.
-This suggests that generating data with an awareness of the characteristics of other classes is more effective.
+Multi-class generation    When generating data under class conditions, a primary consideration is whether to generate data for one class at a time or for all classes simultaneously.
+Single-class generation allows the model to focus on the unique characteristics of a specific class, while multi-class generation enables direct comparison across classes.
+Our experiments indicate that generating data for multiple classes simultaneously in a structured manner (discussed in the following sections) yields samples that more accurately capture the distinctive features of each class.
+This finding suggests that generating data with contextual awareness of other class characteristics enhances the representativeness of synthetic samples.
 Original vs.
-Balanced class ratio    When constructing a prompt to generate multiple classes simultaneously, naive random sampling from the original tabular dataset $T$ often results in an imbalanced ratio of samples for each target class.
-To address this, we consider balancing the sample numbers for minor and major classes rather than adhering to the original data distribution.
-Our findings indicate that balancing the sample numbers significantly improves the quality of generated data for minor classes, demonstrating that training with balanced data helps LLMs learn the true distributions of target classes more effectively.
+Balanced class ratio    When constructing prompts to generate multiple classes simultaneously, simple random sampling from the original dataset $T$ often results in an imbalanced distribution of samples across target classes.
+This imbalance may cause the LLM to overfit to the majority class, limiting its ability to learn the characteristics of the minority class.
+To mitigate this, we employ a balanced sampling approach, equalizing sample numbers for all classes rather than strictly following the original class distribution.
+Our findings reveal that balancing sample sizes substantially improves the quality of generated data for minority classes, indicating that providing balanced data enables LLMs to learn more effectively and replicate the accurate characteristics of all target classes.
 Listing vs.
-Class grouping    Within a prompt, the data examples can either be listed in the order they are sampled without class distinction or grouped by class.
-Empirical evidence suggests that grouping by class leads to a better representation of class-specific feature correlations in the generated data.
-This grouping emphasizes contrasts between different classes and similarities within each group.
-As a result, the LLM generates more distinct samples for each class and more coherent samples within each class.
-Additionally, group-wise prompting can effectively condition LLMs to generate data with specific values.
-Without grouping, it is challenging to constrain LLMs to generate samples with specific attributes, often requiring extensive iterations until the desired samples are produced, which is inefficient.
-However, when examples are provided in groups by class, LLMs learn this and generate samples for each class in groups.
-Thus, grouping is crucial for conditioning the LLM to generate data with specific values and conditions.
-Moreover, grouping can be extended to condition on any arbitrary features other than target classes by grouping samples based on specific requirements.
+Class grouping    Within a prompt, data samples can either be presented sequentially, in the order they are sampled, or grouped by class.
+Grouping emphasizes contrasts between different classes and reinforces similarities within each group, enabling the LLM to generate more distinct and coherent samples for each group.
+Empirical evidence suggests that grouping yields a better representation of class-specific feature correlations in the generated data.
+Class grouping also contributes to constraining the LLM to generate data with specific attributes more efficiently.
+Without grouping, guiding the LLM to produce samples with targeted attributes can be challenging, often requiring numerous iterations to achieve the desired outcome.
+In contrast, grouping samples by specific features facilitates conditioning the LLM to generate data in groups with particular attributes or conditions.
+Although this work primarily focuses on class conditions, this approach can also be applied to feature grouping, enabling controlled generation based on selected criteria.
 2.3 Variable mapping: Original values vs. Unique variable mapping
-As with other LLM-based approaches [4], our approach allows minimal data preprocessing, maintaining the original feature names and values.
-This significantly reduces the need for complex data manipulations often required in existing methods [18, 34] and helps to preserve the integrity of the raw data.
-However, one critical consideration arises when a dataset includes a large number of categorical variables that share the same values (e.g., extensive boolean variables).
-In this case, the data examples in the prompt can become monotonous and repetitive, making it challenging for the model to distinguish between different variables.
-This can significantly degrade the quality of the generated data.
-Moreover, when the LLM receives such input, the number of data samples generated at once can significantly decrease, and the generation efficiency can greatly decline.
-To address this, we propose a unique variable mapping method, which remaps the values of each categorical variable to distinct random alphanumeric strings, as depicted in Fig.
+Unlike existing methods that often require extensive preprocessing and handling of noisy or missing data [19, 40], our approach minimizes these steps, preserving the integrity of the raw data, including original feature names and values, similar to other LLM-based approaches [4].
+However, a critical challenge arises when a dataset contains numerous categorical variables with identical values (e.g., extensive boolean variables).
+In such cases, data examples in the prompt may become monotonous and repetitive, making it difficult for the model to distinguish between variables, which can significantly degrade the quality of the generated data.
+Moreover, when faced with such repetitive input, LLMs tend to struggle to generate valid samples, leading to a notable decline in generation efficiency.
+To address this, we propose a unique variable mapping method, which remaps the values of each categorical variable to distinct random alphanumeric strings, as illustrated in Fig.
 3.
@@
 Although these new values do not hold inherent semantic meaning, they effectively represent categorical distinctions as symbols, allowing the LLM to identify and utilize patterns within them.
-This approach ensures that each variable has a unique representation, making the variables clearly distinguishable and contributing to more accurate and efficient data generation.
+This approach ensures that each variable has a unique representation, making the variables clearly distinguishable and leading to more accurate and efficient data generation.
 2.4 Task specification
-A crucial aspect of the prompt design is to clearly define the task within the prompt, guiding the LLM to generate synthetic data as intended. To achieve this, the following options can be considered:
+A crucial aspect of the prompt design is clearly guiding the LLM to generate synthetic data as intended.
+To achieve this, we consider two options and ultimately find that prompting the LLM to learn the pattern from a structured, consistent format with optional descriptions for variables is more effective for reliable generation than providing explicit instructions.
 Explicit instruction vs.
-Completion triggering    We can clarify the task to the LLM by including explicit instructions to “generate new data samples." However, crafting specific instructions can be challenging, as LLMs are sensitive to minor variations in the prompt, making this approach complex for the tabular data generation task.
-A simpler and more effective method is providing input patterns that LLM can mimic to generate new data, only using data samples without explicit instruction.
-For example, let the $i$-th set of data samples for all $c$ classes be formatted as $\bm{S}_{i}=[\bm{v},\bm{s}_{i,1},\cdots,\bm{s}_{i,c}]$, where $\bm{s}_{i,j}$ consists of $n$ data samples for the $j$-th class.
-Here, the feature name $\bm{v}$ works as a header to indicate the start of a set.
-We input $t$ such sets in a prompt, $[\bm{S}_{1},\bm{S}_{2},\ldots,\bm{S}_{t}]$, creating a pattern and expecting the LLM to generate $\bm{S}_{t\texttt{+}1}=[\bm{v},\bm{s}_{t\texttt{+}1,1},\cdots,\bm{s}_{t\texttt{+}1,c}]$.
-To ensure this, we add a trigger to induce the LLM to complete the sequence.
-That is, we include the header of the new set $[\bm{v}]$ at the end of the prompt, which cues the LLM to generate the rest, $[\bm{s}_{t\texttt{+}1,1},\cdots,\bm{s}_{t\texttt{+}1,c}]$, learning the pattern from the input data that these groups are followed by $\bm{v}$.
-Thus, by creating a pattern in the input prompt, we can condition the LLM to generate a specific number ($c\times n$) of samples for each target class.
-Providing contextual information    Optionally, including contextual information in the prompt can help the LLM understand the data structure and generate accurate samples.
-Therefore, if available, we add line-by-line variable descriptions at the beginning of the prompt.
-Additionally, one can consider clarifying the target class values at the beginning of each group.
-However, we observe that treating the target variable as one of the attributes and placing it as the first variable yields better performance, helping the LLM generate more coherent and accurate samples.
+Completion triggering    To clarify the task for the LLM, a prompt might include explicit instructions, such as “generate new data samples." However, crafting specific instructions can be challenging, as LLMs are highly sensitive to subtle prompt variations, making it inefficient to determine an optimal phrasing for tabular data generation.
+A simpler and more effective approach is to provide patterns within the prompt for the LLM to mimic by formatting data samples without explicit instructions, leveraging its advanced pattern recognition capabilities.
+Formally, let the $i$-th set of data samples across all $c$ classes be structured as $\bm{S}_{i}=[\bm{v},\bm{s}_{i,1},\cdots,\bm{s}_{i,c}]$, where each $\bm{s}_{i,k}$ contains $n$ samples for the $k$-th class.
+Here, $\bm{v}$ represents feature names and serves as a header to indicate the start of a set.
+This approach enables the LLM to recognize the pattern in which each set of data samples is structured and consistently begins with $\bm{v}$.
+We then input $t$ such sets in a prompt, $[\bm{S}_{1},\bm{S}_{2},\ldots,\bm{S}_{t}]$, to establish a recognizable pattern, expecting the LLM to generate the next set, $\bm{S}_{t\texttt{+}1}=[\bm{v},\bm{s}_{t\texttt{+}1,1},\cdots,\bm{s}_{t\texttt{+}1,c}]$.
+To reinforce this, we place a trigger at the end of the prompt by including the new header $[\bm{v}]$ for $\bm{S}_{t\texttt{+}1}$, signaling the LLM to complete the sequence with $[\bm{s}_{t\texttt{+}1,1},\cdots,\bm{s}_{t\texttt{+}1,c}]$.
+By structuring the input prompt in this way, we guide the LLM to generate a consistent number of samples, $n$, for each target class.
+Providing contextual information    Optionally, adding contextual information to the prompt can help the LLM understand the characteristics of the original dataset and generate accurate samples.
+Therefore, when available, we include line-by-line variable descriptions at the beginning of the prompt.
+Additionally, we treat the target class variable as one of the attributes and position it as the first variable to clarify that each group is organized by the target class.
 2.5 Overview of tabular data generation process
-Synthetic data generation starts with creating a prompt template (see Appendix I.5 for examples).
-This template determines the target class and the number of sets, groups, and samples per group to construct demonstrations.
-If numerous categorical variables share the same unique values, we apply unique variable mapping to the dataset.
-Next, we randomly sample data examples from the dataset to fill in this template.
-Based on the provided examples, the LLM generates new data, which is then converted back into a tabular format.
-Instances with categorical variable values not present in the training data are discarded to maintain data authenticity.
-This process is repeated until the desired number of samples is obtained.
-Each iteration involves newly sampled examples from the original dataset, allowing the LLM to encounter various combinations and produce diverse samples.
-Although each individual generation is based on the distribution of the given subset, cumulatively, the iterations enable the model to explore the entire range of the original data.
-This method ensures that the output generated by the LLM comprehensively reflects the original data characteristics (See Appendix G).
-Table 1: Comparison of machine learning classification performance when synthetic data are added to the original data.
-The results are averaged across four classifiers, with each model run five times.
-Complete results, including all baselines and standard deviation values, are provided in Appendix K.
+The synthetic data generation process begins by creating a prompt template and determining whether the unique variable mapping is needed.
+The prompt template specifies parameters such as the target class, sample size per group, the number of groups, and the number of sets to construct demonstrations.
+We then randomly sample data examples from the dataset to populate the template, ensuring that examples do not overlap within each prompt to maintain diversity.
+Using the provided examples, the LLM generates new data in a CSV format, which is subsequently converted into a structured tabular format.
+Instances containing categorical variable values not present in the original dataset are discarded to maintain authenticity.
+This generation process is repeated until the desired sample size is achieved, with new examples sampled with replacement in each iteration.
+This approach enables the LLM to encounter diverse combinations of real samples and produce rich synthetic samples (see Appendix B.1 for analysis).
+Although each individual generation reflects the distribution of the provided subset, successive iterations enable the LLM to cover a broader range of the original data cumulatively.
+This iterative method ensures that the output generated by the LLM comprehensively reflects the characteristics of the original data.
+Final prompting examples are available in Appendix C.5
+3 Experiments
+This section describes the experimental setup and provides a comprehensive evaluation of our method, assessing its effectiveness across diverse real-world datasets.
+Table 1: Comparison of ML classification performance when synthetic data are added to the original dataset.
+Results are averaged across four classifiers, with each model run five times.
+Complete results, including all baselines and standard deviation values, are provided in Table 19.
 Dataset Method #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
@@
 +TVAE [35] +1K 59.78 (+1.66) 72.35 (+1.35) 62.00 (+5.00) 82.69 (-2.31)
++CopulaGAN [22] +1K 21.76 (-36.36) 55.52 (-15.48) 12.80 (-44.20) 98.23 (+13.23)
 +CTAB-GAN+ [40] +1K 54.66 (-3.46) 68.62 (-2.38) 53.00 (-4.00) 84.23 (-0.77)
@@
 +TVAE [35] +1K 87.77 (-0.04) 91.47 (+0.25) 83.37 (+0.54) 99.56 (-0.05)
++CopulaGAN [22] +1K 83.60 (-4.21) 86.61 (-4.61) 73.37 (-9.46) 99.86 (+0.25)
 +CTAB-GAN+ [40] +1K 82.35 (-5.46) 86.28 (-4.94) 72.83 (-10.00) 99.74 (+0.13)
@@
 +TVAE [35] +1K 71.12 (+0.11) 73.25 (+0.04) 68.15 (+0.26) 78.34 (-0.18)
++CopulaGAN [22] +1K 71.23 (+0.22) 73.32 (+0.11) 68.37 (+0.48) 78.26 (-0.26)
 +CTAB-GAN+ [40] +1K 71.03 (+0.02) 73.15 (-0.06) 68.13 (+0.24) 78.17 (-0.35)
@@
 +TVAE [35] +20K 66.96 (+0.06) 76.80 (+0.35) 59.13 (+1.85) 94.48 (-1.13)
++CopulaGAN [22] +20K 66.75 (-0.15) 76.73 (+0.28) 59.16 (+1.88) 94.29 (-1.32)
 +CTAB-GAN+ [40] +20K 66.49 (-0.41) 76.42 (-0.03) 58.14 (+0.86) 94.70 (-0.91)
@@
 +TVAE [35] +10K 54.79 (-0.08) 41.96 (-0.11) 59.96 (-0.04) 60.71 (-0.02)
++CopulaGAN [22] +10K 54.27 (-0.60) 41.59 (-0.48) 59.73 (-0.27) 59.97 (-0.76)
 +CTAB-GAN+ [40] +10K 54.24 (-0.63) 41.52 (-0.55) 59.63 (-0.37) 60.01 (-0.72)
@@
 +TVAE [35] +1K 90.45 (-3.78) 92.20 (-2.88) 86.36 (-4.78) 98.04 (-0.98)
++CopulaGAN [22] +1K 86.73 (-7.50) 88.71 (-6.37) 78.41 (-12.73) 99.02 (0.00)
 +CTAB-GAN+ [40] +1K 27.46 (-66.77) 58.07 (-37.01) 16.14 (-75.00) 100.0 (+0.98)
@@
 +Ours +1K  94.80 (+0.57) 96.39 (+1.31) 95.23 (+4.09) 97.55 (-1.47)
-3 Experiments
-Datasets    We use six real-world classification datasets from diverse domains to evaluate our method: Travel, Sick, HELOC, Income, Diabetes, and Thyroid.
-The Thyroid dataset is released after the training cut-off date for GPT-3.5-turbo-0613, which provides a more rigorous validation of our approach on the completely unseen dataset.
-For binary classification datasets, the minor class is set to be one.
+Summary    We conduct three unique experiments to analyze the utility of our method in enhancing ML classification performance: augmenting the original dataset with generated data (Tables 119, Figs.
+214), augmenting only the minority class similar to SMOTE (Table 6), and using only generated data (Table 7).
+Additionally, we evaluate our approach using three LLMs: Mistral, Llama2, and gpt3.5-turbo (Table 2, Figs.
+45).
+We perform ablations on prompt elements to compare classification performance (Tables 35) and conduct unique analyses of token usage and LLM generation efficiency (Tables 489).
+To examine feature correlations, we separately analyze minority and majority classes, comparing results across all prompt variations (Figs.
+45).
+We further investigate how varying the number of generated samples affects classification performance (Figs.
+67).
+Using a toy dataset, we analyze how different prompts influence the accuracy of generated data distributions (Figs.
+19).
+Lastly, we explore the sampling of input examples and their corresponding outputs in LLMs using the toy dataset, providing insights into the variability and reliability of generated data (Figs.
+1011).
+3.1 Experimental setup
+Datasets    We evaluate our method using six real-world public tabular classification datasets from diverse domains: Travel (Marketing), Sick (Healthcare), HELOC (Finance), Income (Social science), Diabetes (Healthcare), and Thyroid (Healthcare).
+The Thyroid dataset, released after the training cut-off date for GPT-3.5-turbo-0613, provides a more rigorous validation of our approach on completely unseen data.
+For binary classification datasets, the minority class is designated as one.
 All duplicate data samples are removed from the datasets.
-Subsequently, we randomly split them into 80% training and 20% test sets.
-More details for the datasets are provided in Appendix I.2.
-Evaluation measure    We evaluate our generation methodology from three critical aspects: ML classification performance, data generation success rate and stability with LLM in-context learning, and validity of generated data in terms of feature correlation and value range.
-For the classification task, we measure the F1 score, sensitivity, specificity, and balanced accuracy (BAL ACC).
-To measure the feature correlation, we use Pearson correlation and Cramér’s V correlation.
-Baseline models    In our study, we compare our method with various generative models for tabular data generation, including TVAE [30], CopulaGAN [21], CTGAN [30], CTAB-GAN [33], CTAB-GAN+ [34], GReaT [4], and TabDDPM [18].
-This comprehensive evaluation assesses the effectiveness of our approach across different data synthesis paradigms.
-To compare the prompting method, we also use the prompts used in CuratedLLM [25] and LITO [31].
-Experimental details    Our method utilizes the OpenAI GPT-3.5 models, the GPT-3.5-turbo-0613 and GPT-3.5-turbo-16k-0613 versions, Mistral-7b-v0.1 [14], and Llama-2-7b [27].
-Unless stated otherwise, four machine learning classification models are used: XGBoost [8], CatBoost [23], LightGBM [16], and Gradient boosting classifier [11].
-We execute five independent runs for each model and report the average values.
-Additional details are provided in Appendix I.
-Table 2: Comparison of classification performance using synthetic data generated by open-source LLMs, Llama2 and Mistral, with the proposed prompting strategy, utilizing the Gradient Boosting classifier.
-We report the average performances for five runs.
-#syn denotes the number of synthetic samples integrated into the original dataset.
+Each dataset is randomly split into 80% training and 20% test sets.
+We retain the original data, including missing or noisy features.
+The exception is the Sick dataset, where we follow the source’s method.
+Further details are provided in Appendix C.2.
+Evaluation measure    We evaluate our method based on ML classification performance, generation efficiency, feature correlation, and analysis of generated data distribution.
+For classification, we report F1 score, sensitivity, specificity, and balanced accuracy (BAL ACC).
+Feature correlation is measured using Pearson correlation for numerical variables and Cramér’s V correlation for categorical variables.
+Baseline models    In our study, we compare our method with various generative models for tabular data, including SMOTE [5], SMOTENC [5], TVAE [35], CopulaGAN [22], CTGAN [35], CTAB-GAN [39], CTAB-GAN+ [40], GReaT [4], and TabDDPM [19].
+To compare prompting methods, we also use the prompts from CuratedLLM [27] and LITO [37].
+Experimental details    Our method utilizes the GPT-3.5 models (GPT-3.5-turbo-0613 and GPT-3.5-turbo-16k-0613), Mistral-7b-v0.1 [15], and Llama-2-7b [29].
+Unless stated otherwise, we use the GPT-3.5 model for our method.
+The number of synthetic data samples is based on the size of the original datasets.
+Across all experiments, unless stated otherwise, we report results from four top-performing ML classifiers: XGBoost [7], CatBoost [24], LightGBM [17], and Gradient boosting classifier [10], known for their strong performance, often surpassing recent deep learning models on tabular datasets.
+Each classifier is executed in five independent runs, with results averaged over a total of 20 runs to ensure robustness.
+Further details are available in Appendix C.
+Table 2: Comparison of classification performance using synthetic data generated by different LLMs with our method.
+We report the average performances of five runs of gradient boosting classifier.
+#syn denotes the number of synthetic samples added to the original dataset.
 Dataset Method Model #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
@@
 3.2 Machine learning classification performance using the synthetic data
-We assess the quality of the synthetic data samples by comparing the machine learning classification performance averaged over 20 experiments.
+We evaluate the quality of the synthetic data samples by assessing the machine learning classification performance when the synthetic data are added to the original dataset.
+Here, our method utilizes the GPT-3.5-turbo model.
 As shown in Table 1, our method achieves state-of-the-art F1 scores and balanced accuracy across all six datasets, surpassing the baselines that require model training.
 Notably, our method is the only one that consistently outperforms the original data in both F1 score and balanced accuracy across all six datasets.
-Other methods often result in worse performance than the original data, indicating that the data generated by baseline methods disrupts the original data distribution.
-These findings underscore the robustness and effectiveness of our method, affirming its superiority over baseline models, even for challenging datasets.
-Moreover, baseline methods such as GReaT and TabDDPM exhibit significantly lower sensitivity compared to specificity, particularly for Travel and Income.
+For a machine learning model to perform well in classification, the correlation between input data and labels in the training data must be precise.
+Thus, the consistent improvement across six datasets demonstrates that the data generated by our method aligns well with the class labels.
+Other methods often result in worse performance than the original data, indicating that the data generated by baselines disrupts the original data distribution.
+These findings underscore the robustness and effectiveness of our method, affirming its superiority over baselines, even for challenging datasets.
+Moreover, baselines such as GReaT and TabDDPM exhibit significantly lower sensitivity compared to specificity, particularly for Travel and Income.
 Their high balanced accuracy is due to high specificity, but their severely low sensitivity indicates a failure to generate appropriate minority class data.
@@
 Overall, our findings demonstrate that our approach is generally applicable across various imbalanced tabular datasets and excels in generating high-quality samples that improve ML classification performance.
-For detailed confusion matrix results, see Appendix J, and for further analyses exploring the impact of adding synthetic data only to the minor class or using synthetic data alone, see Appendix A.
+Further analyses exploring the impact of augmenting the original dataset for the minority class (Table 6) and replacing the original dataset with synthetic data (Table 7) are available in Appendix A.
 3.3 Open-source LLMs
@@
 CSV-style multi 91.95$\pm$0.00 - -
-Table 4: Comparison of token efficiency and generation stability on the Income dataset.
+Table 4: Comparison of token usage and generation efficiency across ablated methods on the Income dataset.
+Results are based on 100 inferences, each with 20 random input samples.
+Input tokens indicates the number of tokens required in the LLM prompt for a fixed number of input samples.
+Output samples shows the average number of synthetic samples generated per iteration.
+Success rate measures the ratio of inferences that generate at least one valid data sample.
 Format Class Balance Group Unique #set Input tokens $\downarrow$ Output samples $\uparrow$ Success rate $\uparrow$
@@
 CSV-style multi 2 1014.6 10.52 94%
-Table 5: Comparison of classification performance on task specification elements in prompt design, averaged across four classifiers: XGBoost, CatBoost, LightGBM, and Gradient Boosting classifier.
-Methods marked with an asterisk (*) use prompts as proposed in the respective papers.
+Table 5: Comparison of classification performance on the Sick dataset for task specification elements in prompt design.
+Results are averaged across XGBoost, CatBoost, LightGBM, and gradient boosting classifier.
+Methods marked with an asterisk (*) use the prompt designs proposed in the respective papers.
+#syn denotes the number of synthetic samples added to the original dataset.
 Method #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
@@
 Ours +1K 88.71$\pm$1.98 92.93$\pm$0.91 86.41$\pm$1.85 99.44$\pm$0.27
-Figure 4: The difference between Cramér’s V correlation matrices of real and synthetic datasets for categorical variables in the Travel dataset.
-More intense colors indicate larger differences, with positive differences shown in red and negative differences shown in blue.
-Black indicates where the correlation is not measured since only one unique value is generated for those variables.
 3.4 Exploring optimal prompt design through ablation studies
-We investigate key prompt design elements focusing on ML performance and generation efficiency, as detailed in Tables 3 and Table 4.
-Token efficiency is measured by the number of input tokens required for a fixed number of input samples.
-LLM generation efficiency is evaluated based on the average number of samples generated per iteration and the success rate of generating at least one valid data sample per inference.
-This evaluation is based on 100 iterations, each with 20 random input samples.
+We investigate key prompt design choices and validate our method through extensive ablation studies across multiple datasets from diverse domains, focusing on ML performance and generation efficiency, as detailed in Tables 3 and 4, respectively.
+Given a fixed number of input samples, we evaluated (1) the number of input tokens required, (2) the number of valid generated samples, and (3) the generation success rate.
 Our results indicate that CSV-style prompting generally outperforms sentence-style prompting in ML performance with the same number of input samples.
@@
 Repeating the set twice within the prompt further improves generation efficiency, boosting the average output to over 10 samples and the success rate to 94%, by creating input patterns that LLMs can mimic to generate data.
-Unique variable mapping also significantly enhances LLM generation efficiency and success rates, as detailed in Appendix B.
+Unique variable mapping also significantly enhances LLM generation efficiency and success rates, as detailed in Table 8 of Appendix A.3.
+Figure 4: Difference between Cramér’s V correlation matrices of real and synthetic datasets for categorical variables in the Travel dataset.
+More intense colors indicate larger differences, with positive differences shown in red and negative differences shown in blue.
+Black indicates where the correlation is not measured since only one unique value is generated for those variables.
 Furthermore, we conduct an ablation study on the task specification elements, as shown in Table 5.
@@
 Providing variable descriptions improves balanced accuracy but reduces the F1 score, suggesting that this element may be optional rather than a definitive improvement.
-For a comparison of generation efficiency using these methods, see Appendix B.
-3.4 Analysis on feature correlation
+A comparison of generation efficiency using these methods is provided in Table 9 of Appendix A.3.
+3.5 Analysis of feature correlation
 We compare the feature correlations of the categorical variables between original and synthetic data, as illustrated in Fig.
 4.
-Our method, utilizing Mistral and Llama2, produces data that most closely mirrors the original feature correlations for both minor and major classes.
-Moreover, when compared to the ablated versions, our grouping method significantly contributes to preserving feature correlation for the minority class.
-For additional feature correlation results on the Sick dataset, see Appendix C.
+Our synthetic data generated with Mistral and Llama2 exhibit the closest feature correlation with the original data for both minority and majority classes, outperforming the baselines and ablated versions.
+Additionally, our grouping method significantly contributes to preserving feature correlation for the minority class.
+These findings indicate that our method generates data that more accurately matches each class, particularly for underrepresented classes, compared to other methods.
+Additional results on the Sick dataset are available in Fig.
+5 of Appendix A.3.
+4 Limitations and future work
+When the training dataset is large and cannot be fully included in the LLM prompt due to token size limitations, only a subset of the data can be used as examples for generating samples.
+If these prompt samples do not fully represent the original data distribution, the generated data may be incomplete and of low quality, as LLMs are limited to producing data based only on the patterns present in the input.
+To overcome this, our method employs multiple rounds of random sampling with replacement to create a synthetic dataset that more comprehensively represents the original distribution, resulting in improved machine learning classification performance (a comprehensive analysis is provided in Appendix B.1).
+However, this approach still carries the risk that samples may not fully capture the original data distribution.
+Future research will focus on developing methods to identify key examples that more accurately represent the entire dataset.
 5 Conclusion
-This study demonstrates the effectiveness of using LLMs for tabular data generation and introduces an optimal solution for prompt design to generate realistic and reliable data without additional training, particularly addressing class imbalance.
-Our extensive evaluations demonstrate the robustness and effectiveness of our method, which significantly improves ML classification performance on the six real-world tabular datasets.
-This highlights its practicality and potential impact on improving model performance in various application scenarios, contributing significantly to the field.
+This study demonstrates the effectiveness of using LLMs for synthetic tabular data generation.
+We introduce EPIC, a simple yet effective solution for generating realistic, high-quality data without additional training, specifically designed to address class imbalance.
+Our method demonstrates significant improvements over state-of-the-art generation models across six real-world public datasets, generating data with highly accurate feature correlations and significantly improving ML classification accuracy for minority classes.
+These results underscore the substantial impact of our approach in real-world applications, making a significant contribution to the field of tabular data research.
+Acknowledgements
+This work was supported by Institute for Information & communications Technology Promotion(IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II190075 Artificial Intelligence Graduate School Program(KAIST)), Electronics and Telecommunications Research Institute(ETRI) grant funded by the Korean government (24ZB1220, Fundamental Technology Research for Human-Centric Autonomous Intelligent Systems), and the National Supercomputing Center with supercomputing resources including technical support (KSC-2024-CRE-0028).
 Appendix
 This supplementary material enhances the main manuscript by providing detailed experimental results and additional visualizations.
-  • Appendix A presents more experimental results on the machine learning classification for imbalanced classes.
-  • Appendix B offers complete additional ablation study results.
-  • Appendix C provides feature correlation results for the Sick dataset.
-  • Appendix D studies the impact of sample size on the machine learning performance.
-  • Appendix E comprehensively analyzes the experiments conducted on the toy dataset.
-  • Appendix F discusses about the broader impacts of our research.
-  • Appendix G outlines limitations of our work.
-  • Section H provides related work.
-  • Appendix I provides comprehensive information on the datasets, comparison baselines, implementation details, and prompt design and examples.
-  • Appendix J presents confusion matrix results, highlighting the effectiveness of our method in addressing minor class problems.
-  • Appendix K shows the complete results of Table 1.
+  • Appendix A provides additional experimental results on the machine learning classification for imbalanced data, ablation study, and feature correlation analysis.
+  • Appendix B studies the impact of sample size on the machine learning performance and comprehensively analyzes the experiments conducted on the toy dataset.
+  • Appendix C provides comprehensive information on the datasets, baselines, implementation details, and prompt design and examples.
+  • Appendix D provides related work.
+  • Appendix E discusses the broader impacts of our research.
+  • Appendix F presents confusion matrix results and the complete results of Table 1.
 Appendix A Additional experimental results
-Our experiments with synthetic tabular data for imbalanced classes follow three main approaches: (1) Adding synthetic data to both minor and major classes of the existing dataset and evaluating machine learning classification performance.
+Our experiments with synthetic tabular data for imbalanced classes follow three main approaches: (1) Adding synthetic data to both minority and majority classes of the existing dataset and evaluating machine learning classification performance.
 The results of this approach are presented in Table 1 of the main manuscript.
-(2) Adding synthetic data only to the minor class, with detailed results provided in Appendix A.1.
+(2) Adding synthetic data only to the minority class, with detailed results provided in Appendix A.1.
 (3) Evaluating classification performance using only synthetic data without the original dataset, as discussed in Appendix A.2.
-Table 6: Comparison of classification performance for imbalanced data by generating synthetic data for the minority class to match the size of the majority class.
-The average performances of the Gradient Boosting Classifier over five runs are reported.
+Moreover, Appendix A.3 presents further ablation study results, and Appendix A.4 provides a feature correlation analysis for the sick dataset.
+Table 6: Comparison of binary classification performance when augmenting the minority class with synthetic data to balance class sizes.
+Average performance of the gradient boosting classifier is reported over five runs.
+#syn denotes the number of synthetic samples added to the original dataset.
 Dataset Method #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
@@
 +Ours +13,487 71.09$\pm$0.02 83.60$\pm$0.01 83.53$\pm$0.03 83.66$\pm$0.00
-A.1 Comparative analysis with adding the new dataset only to the minor class
-One of the major advantages of our method is its ability to generate high-quality synthetic data for minor classes, even with limited samples.
-Similar to oversampling techniques like SMOTE, we add synthetic dataset only to the minor class and evaluate machine learning classification performance.
-To balance the ratio of major to minor classes, we add only minor class samples to the original data, then measure the classification performance of the machine learning model.
+A.1 Comparative analysis of augmenting the original dataset for the minority class
+One of the major advantages of our method is its ability to generate high-quality synthetic data for minority classes, even with limited samples.
+Similar to oversampling techniques like SMOTE, we add synthetic dataset only to the minority class and evaluate machine learning classification performance.
+To balance the ratio of majority to minority classes, we add only minority class samples to the original data and then measure the classification performance of the machine learning model.
 For this task, we additionally investigate the effectiveness of our method compared to SMOTE and SMOTENC [5].
-As shown in Table 6, adding minor class samples generated from our method leads to higher F1 scores and balanced accuracy on both Travel and Income datasets, compared to when samples generated by other baselines are added.
-In the Travel dataset, where the number of minor class samples is substantially low, baseline models, especially, CopulaGAN and CTGAN, demonstrate markedly low sensitivity, indicating their failure to accurately learn the minor class data distribution.
-In contrast, our method demonstrates its efficacy in producing high-quality data for minor classes on the Travel dataset.
+As shown in Table 6, adding minority class samples generated from our method leads to higher F1 scores and balanced accuracy on both Travel and Income datasets, compared to when samples generated by other baselines are added.
+In the Travel dataset, where the number of minority class samples is substantially low, baselines, especially CopulaGAN and CTGAN, demonstrate markedly low sensitivity, indicating their failure to accurately learn the minority class data distribution.
+In contrast, our method demonstrates its efficacy in producing high-quality data for minority classes on the Travel dataset.
 Compared to the Original, it significantly improves the F1 score by 12.13%p, underscoring its superior performance of our approach.
-A.2 Comparative analysis of classification performance using only synthetic data
-Table 7: Comparison of classification performance using only synthetic data for training. We report average values for five runs employing Gradient Boosting Classifier.
+A.2 Comparative analysis of replacing the original dataset with synthetic data
+Table 7: Comparison of classification performance using only synthetic data for training. Average performance of the gradient boosting classifier is reported over five runs.
 Dataset Method #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
@@
 We compare the classification performance using only synthetic data for training, as shown in Table 7.
-For both the Travel and HELOC datasets, the classification model trained on synthetic datasets generated by our model outperform the one trained on original data in terms of F1 score and Sensitivity.
+For both the Travel and HELOC datasets, the classification model trained on synthetic datasets generated by our model outperforms the one trained on original data in terms of F1 score and sensitivity.
 This improvement in performance suggests that the data generated by our method accurately represents the distribution of the original data.
 Furthermore, it highlights the effectiveness of our class-balanced generation approach in improving the performance of classification models.
-Table 8: Comparison of token efficiency and generation stability for ablating prompt design elements on the Sick dataset
+A.3 Additional ablation study results
+This section presents additional ablation study results on the Sick datasets.
+We evaluate the token efficiency and generation stability of ablated versions based on 100 inferences, each with 20 random input samples.
+As shown in Table 8, for the Sick dataset, due to its repetitive, monotonous values, the number of generated samples is low, even with grouping.
+Using a unique variable mapper in such cases significantly increases the number of samples generated per attempt from 6.93 to 17.68.
+This improvement demonstrates that our approach effectively controls the quantity of generated data samples, ensuring the stable production of synthetic tabular data.
+Moreover, the ability to generate multiple data samples in a single inference can significantly lower generation time and costs.
+We also compare the efficiency and stability of generation when ablating task specification elements in our prompt design, as shown in Table 9.
+The prompt design from Curated LLM [27], despite requiring the highest number of input tokens, achieves high performance in terms of output samples and success rate.
+However, the generated data shows relatively low quality, resulting in a low classification performance with an F1 score and balanced accuracy, as discussed in Section 3.4.
+In contrast, the prompt design of LITO [37] is the most efficient in terms of input tokens but produces the fewest output samples and has the lowest success rate, indicating instability in synthetic data generation.
+Our proposed method balances efficiency and performance, using approximately half the input tokens compared to CuratedLLM’s prompt while generating a high number of output samples and maintaining a high success rate.
+Additionally, the synthetic data generated by our method achieves the highest classification performance in terms of F1 score, as shown in Section 3.4.
+When class distinctions are provided, the average number of output samples decreased from 17.68 to 9.70.
+This result indicates that it is more robust to distinguish groups with simple numbering, such as A and B, rather than explicitly specifying class information in each group.
+Including variable descriptions slightly reduces the number of output samples and success rate compared to not including them, but the difference is not significant.
+This suggests that the grouping strategy plays a more crucial role in robust data generation than the presence of variable descriptions.
+Table 8: Comparison of token usage and generation efficiency across ablated methods on the Sick dataset.
+Results are based on 100 inferences, each with 20 random input samples.
+Input tokens indicates the number of tokens required in the LLM prompt for a fixed number of input samples.
+Output samples shows the average number of synthetic samples generated per iteration.
+Success rate measures the ratio of inferences that generate at least one valid data sample.
 Format Class Balance Group Unique #set Input tokens $\downarrow$ Output samples $\uparrow$ Success rate $\uparrow$
@@
 CSV-style multi 2 2060.8 17.68 95%
-Table 9: Comparison of token efficiency and generation stability for task specification in prompt design on the Sick dataset.
+Table 9: Comparison of token usage and generation efficiency for task specification ablation on the Sick dataset. Results are based on 100 inferences, each with 20 random input samples.
 Method Input tokens $\downarrow$ Output samples $\uparrow$ Success rate $\uparrow$
@@
 Ours 2060.8 17.68 95%
-Appendix B Additional ablation study results
-This section presents additional ablation study results on the Sick datasets.
-We evaluate the token efficiency and generation stability of ablated versions based on 100 iterations, each with 20 random input samples.
-As shown in Table 8, for the Sick dataset, due to its repetitive monotonous values, the number of generated samples is low, even with grouping.
-Using a unique variable mapper in such cases significantly increases the number of samples generated per attempt from 6.93 to 17.68.
-This improvement demonstrates that our approach effectively controls the quantity of generated data samples, ensuring the stable production of synthetic tabular data.
-Moreover, the ability to generate multiple data samples in a single inference significantly can lowers generation time and costs.
-We also compare the efficiency and stability of generation when ablating task specification elements in our prompt design, as shown in Table 9.
-The prompt design from CuratedLLM [25], despite requiring the highest number of input tokens, achieves high performance in terms of output samples and success rate.
-However, the generated data shows relatively low quality, resulting in a low classification performance with an F1 score and balanced accuracy, as discussed in Section 3.3.
-In contrast, the prompt design of LITO [31] is the most efficient in terms of input tokens but produces the fewest output samples and has the lowest success rate, indicating instability in synthetic data generation.
-Our proposed method balances efficiency and performance, using approximately half the input tokens compared to CuratedLLM’s prompt while generating a high number of output samples and maintaining a high success rate.
-Additionally, the synthetic data generated by our method achieves the highest classification performance in terms of F1 score, as shown in Section 3.3.
-When class distinctions are provided, the average number of output samples decreased from 17.68 to 9.70.
-This result indicates that it is more robust to distinguish groups with simple numbering, such as A and B, rather than explicitly specifying class information in each group.
-Including variable descriptions slightly reduces the number of output samples and success rate compared to not including them, but the difference is not significant.
-This suggests that the grouping strategy plays a more crucial role in robust data generation than the presence of variable descriptions.
-Figure 5: The differences between Pearson correlation matrices of real data and synthetic data for numerical variables in the Sick dataset.
+Figure 5: Difference between Pearson correlation matrices of real and synthetic data for numerical variables in the Sick dataset.
 More intense colors indicate larger differences, with positive differences shown in red and negative differences shown in blue.
-Appendix C Analysis of feature correlation between numerical variables
+Black indicates where the correlation is not measured since only one unique value is generated for those variables.
+A.4 Analysis of feature correlation between numerical variables
 In the main manuscript, we visualize the feature correlation between categorical variables.
@@
 5.
-As a result, our methodology consistently shows feature correlations similar to the original data for both minor and major classes.
-In contrast, other baselines either show significant deviations in feature correlation for all classes or perform well for major classes but not for minor ones.
-These results demonstrate that our methodology shows superior performance compared to the baselines in generating high-quality data that accurately represents both minor and major classes within the original dataset.
-Appendix D Impacts of sample size on the classification performance
-Figure 6: Impact of sample size on the classification performance on the Income and Diabetes datasets when classifiers are trained exclusively on synthetic samples generated from our proposed method.
+As a result, our methodology consistently shows feature correlations similar to the original data for both minority and majority classes.
+In contrast, other baselines either show significant deviations in feature correlation for all classes or perform well for majority classes but not for minority ones.
+These results demonstrate that our methodology shows superior performance compared to the baselines in generating high-quality data that accurately represents both minority and majority classes within the original dataset.
+Appendix B Additional analysis of EPIC
+This section provides an in-depth analysis of the proposed method, EPIC.
+Appendix B.1 analyzes the effect of sample size on classification performance.
+Appendix B.2 discusses the significance of the observed performance improvements.
+Appendix B.3 provides complete qualitative results on the toy dataset.
+Finally, Appendix B.4 explores tabular data classification with in-context learning methods.
+B.1 Impacts of sample size on the classification performance
+Figure 6: Classification performance on the Income dataset when synthetic data generated by our proposed method are added to the original dataset.
+We experiment with varying synthetic sample sizes, comparing data sampling methods: with replacement (w/ rep.) and without replacement (w/o rep.).
+EPIC samples with replacement.
+GPT-3.5-turbo is used for the experiment.
+Figure 7: Classification performance on the Income and Diabetes datasets using only synthetic data generated by our proposed method for training.
 Black dashed lines denote the F1 score using only the original data.
-Figure 7: Complete generation results on an imbalanced toy dataset.
-Major class is shown in blue and minor class in red.
-Our approach via in-context learning of LLM achieves (1) clear class boundaries, (2) accurate feature correlations, (3) proper value ranges, (4) well-represented relationships between categorical (last row) and numerical variables, and (5) better diversity compared to the fine-tuned GReaT [4] model.
-We investigate how varying the number of synthetic samples affects classification performance.
+GPT-3.5-turbo is used for our method.
+Adding synthetic data to the original dataset    Data samples in prompts can be selected either with or without replacement.
+Our method uses sampling with replacement to generate a large and diverse dataset, as sampling without replacement limits the amount of synthetic data to the number of actual data points.
+While we ensure that there are no overlapping examples within each prompt to maintain diversity, each prompt is constructed using sampling with replacement.
+We conduct experiments to analyze (1) how much datasets can be enlarged and (2) how this impacts classification performance by comparing sampling with and without replacement.
+As shown in Fig.
+6, performance improves steadily in both scenarios as the volume of generated data increases.
+However, the improvement is constrained when sampling without replacement due to the limited number of possible samples.
+When sampling with replacement, as the dataset size expands, there is a noticeable improvement in the balance between sensitivity and specificity, which contributes to enhanced overall performance, including gains in balanced accuracy and F1 score.
+Generating up to 40K synthetic data points resulted in even better performance than the 20K synthetic data points reported in Table 1 of our main manuscript.
+We also observed that as the volume of generated data continues to increase, the gains in balanced accuracy and F1 score eventually plateau, indicating diminishing returns and suggesting that further data generation beyond a certain point offers limited additional benefit.
+Using only synthetic data    We investigate how varying the number of synthetic samples affects classification performance.
 This analysis is conducted on the Income and Diabetes datasets, and the results are shown in Fig.
-6.
+7.
 For the Income dataset, sample sizes range from 500 to 10,000, while for Diabetes, the range was 500 to 20,000.
-For both datasets, a consistent increase in the F1 score is observed as the number of samples grow.
+For both datasets, a consistent increase in the F1 score is observed as the number of samples grows.
 This result demonstrates that our method can adequately represent the original data distribution by generating a sufficient number of synthetic data samples.
-Figure 8: The illustration of the input samples (blue and orange) and the generated samples (green and red) for each inference on the imbalanced toy dataset when the our proposed approach received 180 input samples.
-Empty spaces in the visualization indicate instances where the model failed to generate samples.
-To produce a total of 1000 samples, the model performed 23 inferences.
-Figure 9: The illustration of the input samples (blue and orange) and the generated samples (green and red) for each inference on the imbalanced toy dataset when the ablated model without grouping received 180 input samples.
-Empty spaces in the visualization indicate instances where the model failed to generate samples.
-To produce a total of 1,000 samples, the model performed 37 inferences.
-Appendix E Complete qualitative results on toy dataset
+Figure 8: Average classification results across six datasets compared with the closest baselines.
+Table 10: Performance improvement after adding synthetic data to the original dataset. Results are reorganized from Table 1 of our main manuscript.
+Method Dataset Improvement from the original data (%p)
+F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
+GReaT Travel 2.83 1.86 1.8 1.92
+Sick -0.58 -0.39 -0.76 -0.01
+HELOC -0.66 -0.25 -1.67 1.18
+Income 1.05 1.06 3.41 -0.31
+Diabetes -0.09 -0.09 -0.02 -0.12
+Thyroid -2.92 -2.62 -5.23 0.00
+TabDDPM Travel -4.92 -3.3 -6.6 0.00
+Sick -2.64 -1.92 -3.81 -0.04
+HELOC -0.36 -0.32 -0.38 -0.26
+Income -0.05 0.05 0.42 -0.31
+Diabetes -0.23 -0.24 -0.09 -0.18
+Thyroid 0.16 1.18 4.31 -1.96
+Ours Travel 8.53 7.23 21 -6.54
+Sick 0.9 1.71 3.58 -0.17
+HELOC 0.91 0.45 2.07 -1.17
+Income 2.26 2.7 9.17 -3.76
+Diabetes 0.07 0.07 0.04 0.09
+Thyroid 0.57 1.31 4.09 -1.47
+B.2 Significance of performance improvements
+We analyze the significance of performance improvements achieved by EPIC.
+As shown in Fig.
+6, while the gains in F1 score and balanced accuracy between 0 and 40K samples may seem modest, the sensitivity actually increases significantly from around 50% to 70%.
+This improvement balances performance across all classes, leading to a more meaningful outcome and greatly enhancing model usability.
+Baselines often learn biases in the training data, resulting in abnormally high specificity at the expense of sensitivity.
+This imbalance can inflate balanced accuracy or F1 scores, but such performance is ineffective for real-world tasks.
+In contrast, our method significantly enhances sensitivity with only a slight reduction in specificity.
+As a result, our method achieves the most balanced performance across all four metrics, as shown in Fig.
+8.
+Our method is the only approach among the seven baselines that consistently improves both F1 score and balanced accuracy compared to using only the original data (Table 19).
+Extensive experiments across six real-world datasets from diverse domains (finance, healthcare, marketing, and social science) using four classifiers, each tested five times, demonstrate the state-of-the-art performance of our method, with significant gains in the highly imbalanced Travel, Sick, Income, and Thyroid datasets (Fig.
+13).
+While GReaT and TabDDPM are the closest baselines, they exhibit inconsistent performance.
+As highlighted in Table 10, they reduce original data performance in more than half of the cases (red).
+For example, both underperform on HELOC and Diabetes.
+In stark contrast, our method consistently outperforms the original data across all datasets (blue).
+In the challenging Diabetes dataset, only our method improves over the original data.
+While specificity decreases in all methods, our method still achieves a superior balance of metrics, leading to greater practicality.
+These results underscore that our method offers the greatest practical utility and a meaningful performance advantage among the models tested.
+Figure 9: Complete generation results on an imbalanced toy dataset with majority and minority classes.
+Our approach, leveraging in-context learning with LLMs, achieves (1) distinct class boundaries, (2) accurate feature correlations, (3) well-matched value ranges, (4) robust numerical-categorical relationships (last row), and (5) comprehensive data distribution coverage, with improvements over its ablated versions and the fine-tuned GReaT model [4].
+B.3 Complete results on toy dataset
 In Fig.
-7, we illustrate the complete qualitative results on the toy dataset.
+9, we illustrate the complete qualitative results on the toy dataset.
 We assume a challenging multivariate data scenario in an imbalanced binary classification situation, where variables with and without correlations are mixed.
@@
 These variables have strong correlations in pairs.
-There are 11 numerical variables and 1 categorical variable, and the categorical variable is shown on the x-axis in the last row of Fig.
-7.
-In the figure, the major class is shown in blue, and the minor class is shown in red.
-The generated results can be evaluated from four aspects.
-Firstly, whether each data sample is correctly generated to belong to each class, i.e., whether the class boundary is clear.
-Secondly, whether the shapes formed by the correlation between the variables, such as circles or moons, are well represented.
-Thirdly, whether the values generated fall within the acceptable range for each variable.
-Finally, whether the relationship between the categorical and numerical variables is well represented.
+There are 11 numerical variables and one categorical variable, and the categorical variable is shown on the x-axis in the last row of Fig.
+9.
+In the figure, the majority class is shown in blue, and the minority class is shown in red.
+The generated results are evaluated from four aspects: (1) assessing whether generated data samples correctly belong to their respective classes, ensuring clear class boundaries; (2) examining whether the shapes formed by variable correlations, such as circles or moon patterns, are accurately represented; (3) evaluating whether the generated values fall within an acceptable range for each variable; and (4) analyzing whether relationships between categorical and numerical variables are accurately captured.
 The baseline method, GReaT [4], has successfully learned from the training data to generate class-separated samples.
@@
 Ours, despite generating samples without training, shows similar or better quality.
+Figure 10: Illustration of the input samples (majority and minority) and the corresponding generated samples (majority and minority) for each inference of EPIC on the imbalanced toy dataset.
+In this example, with 180 input samples per inference, EPIC requires 23 inferences to generate 1,000 samples.
+Empty boxes represent cases where the model fails to generate valid samples.
+Figure 11: Illustration of the input samples (majority and minority) and the corresponding generated samples (majority and minority) for each inference of the ablated version of our method on the imbalanced toy dataset.
+In this example, with 180 input samples per inference, the ablated method without grouping requires 37 inferences to generate 1,000 samples.
+Empty boxes represent cases where the model fails to generate valid samples.
 Overall, CSV-style generation produces better class separation boundaries compared to sentence-style generation.
-Regarding class presentation, giving a single class can sometimes result in good generation outcomes, but it fails to accurately capture the correlation with the major class, resulting in shapes like squares instead of circles, and it also fails to accurately generate the categorical variable.
-On the other hand, when generating multi-class samples, if the sample numbers are not balanced, the generation results for the minor class can be compromised.
+Regarding class presentation, giving a single class can sometimes result in good generation outcomes, but it fails to accurately capture the correlation with the majority class, resulting in shapes like squares instead of circles, and it also fails to accurately generate the categorical variable.
+On the other hand, when generating multi-class samples, if the sample numbers are not balanced, the generation results for the minority class can be compromised.
 Even when balanced, providing grouping leads to clearer class distinctions.
@@
 10 and Fig.
-9.
+11.
 For each inference, 180 class-balanced samples are input into the LLM prompt.
 To generate a total of 1,000 samples, our method required 23 inferences, whereas the ablated model needed 37 inferences, demonstrating the stability of our method in generating synthetic data samples.
-In terms of data generation quality, our method successfully generates both major and minor classes in most inferences, with the distribution of the generated samples closely mirroring that of the actual input data.
+In terms of data generation quality, our method successfully generates both majority and minority classes in most inferences, with the distribution of the generated samples closely mirroring that of the actual input data.
 In contrast, the ablated model exhibits a higher number of inferences where the generated data deviates from the input data distribution.
-Appendix F Broader Impact
-Our method is generally applicable to any tabular classification task with minimal pre-processing, making it accessible for various research and industrial applications.
-The potential positive impacts include enhanced data accessibility and privacy preservation, which can improve decision-making processes in fields like healthcare and finance.
-Additionally, it reduces the need for using real data, thereby protecting individual privacy.
-However, there are potential negative impacts, such as the risk of data misuse.
-Synthetic data generation could be misused for malicious purposes, such as creating fake datasets to deceive systems or individuals.
-Figure 10: The illustration of the input (blue and orange) and generated (green and red) samples from our method where input samples are exclusively taken from either the left or right side of the circle.
-Appendix G Limitations
-If the input samples do not adequately represent the data, the LLM may fail to generate samples that follow the original data distribution, leading to degraded performance.
-As shown in Fig.
-10, if the input samples do not fully capture the original data shape, the generated data may fail to accurately follow this shape.
-A limitation of our work is that our current method relies on providing a large number of data samples to ensure sufficient representation in in-context learning.
-Future work will focus on increasing efficiency by selectively choosing key samples that can help the LLM generate data effectively with fewer examples.
-Appendix H Related work
-In the field of tabular data synthesis, traditional methods have predominantly focused on sampling techniques, such as SMOTE [6], which have been foundational in addressing class imbalances in tabular data.
-Although useful, these methods often fall short in capturing complex inter-feature correlations [2] and typically lack the versatility needed for regression tasks or addressing multiple feature imbalances.
-Recently, deep-learning-based generative models have demonstrated their efficacy in generating realistic tabular data [7, 9, 15, 30, 33, 34].
-For example, CTAB-GAN [33] introduced a generator that mitigates the problem caused by imbalanced discrete variables.
-Additionally, diffusion models have also exhibited promising results in tabular data generation.
-TabDDPM [18] introduced a diffusion process for both numerical and categorical features in tabular data for accurate data generation.
-Large language model (LLM)-based methods have emerged as a recent trend in tabular dataset generation.
-GReaT [4] surpassed previous methods by fine-tuning LLMs to generate tabular data formatted in natural text.
-However, this fine-tuning process is resource-intensive, requiring extensive training of large models for each dataset.
-As an alternative, leveraging the in-context learning capabilities of LLMs presents a promising solution.
-LLMs have demonstrated the ability to solve various problems, learn and mimic patterns from given input data, and generate new data without additional training for parameter updates [5, 13, 17, 19, 28, 29, 32].
-Despite this potential, there is limited research on what LLMs generate for synthetic tabular data and what the optimal prompts for high-quality tabular data generation would be.
-Previous studies [31, 25] have shown the potential of LLMs for in-context learning of tabular data generation.
-However, these studies have not sufficiently explored optimal prompts for synthetic tabular data generation, leading to instability in the generated samples.
-This instability necessitates additional validation modules and a two-stage structure.
-Given that the performance of in-context learning with LLMs heavily depends on prompt design [10, 26, 28, 32], carefully designing prompts specifically for imbalanced tabular data is essential to maximize LLM capabilities in effectively generating realistic data.
-Recognizing the importance of optimal prompt design and the potential instability in generated data, we propose a novel method for generating high-quality synthetic tabular data through tailored prompting.
-The proposed method is generally applicable to various types of tabular datasets without the need for training, thereby offering a cost-effective solution in data synthesis.
-Appendix I Additional experimental details
-This section provides the illustration of ablated versions (Appendix I.1), comprehensive information on the datasets (Appendix I.2) and comparison baselines (Appendix I.3), implementation details (Appendix I.4), and final prompt design and examples (Appendix I.5).
-I.1 Ablation prompt examples
+B.4 Comparison of advanced tabular classification models
+The ability of LLMs to generate high-quality synthetic data for imbalanced classes in our approach highlights their potential to directly address classification tasks, offering a promising avenue for future research.
+However, the challenge lies in designing effective prompts to fully leverage this potential.
+The performance of LLMs is highly dependent on prompt design, and numerous studies have shown that small variations in prompt crafting can lead to significant differences in outcomes [18, 25].
+Our work shares this motivation, focusing on designing prompts that enable LLMs to effectively generate high-quality tabular data, particularly to address class imbalance.
+While exploring LLMs for direct classification could enhance performance, this direction is beyond the scope of our current study.
+We conduct experiments using recent in-context learning-based tabular classification methods, such as TabPFN [14], which employs a pretrained transformer, and T-Table [33], which leverages LLMs.
+We also test TabR [12], an advanced deep-learning tabular classification model, using its official code.
+As detailed in Table 11, we evaluate classification performance on the Travel dataset using the original and synthetic data.
+For T-Table, we use the GPT-3.5-turbo model, incorporating all original data but limiting synthetic samples to 200 per input prompt due to token constraints.
+To further enhance performance, we employ voting across five inferences [31].
+Table 11: Comparison of F1 scores using robust tabular classification models on the Travel dataset. GB refers to gradient boosting classifier.
+Model Original +Ours +TabDDPM +GReaT
+XGBoost 55.32 67.74 51.06 64.00
+LightGBM 60.00 64.29 55.32 62.50
+CatBoost 57.14 64.41 48.09 57.73
+GB 60.00 70.18 58.33 59.57
+TabPFN 56.00 59.26 55.32 53.06
+T-Table 21.62 30.43 10.53 17.65
+TabR 46.41 60.78 44.88 32.41
+The results indicate that, while the newly introduced classifiers do not outperform traditional ones, adding synthetic data generated by our method to the original data consistently leads to significant performance improvements across all classifiers.
+In contrast, using TabDDPM and GReaT often results in decreased performance.
+Our method uniquely and consistently enhances the performance of different classifiers, demonstrating superior label-matching quality.
+These findings underscore the value of high-quality synthetic data in enhancing classifier performance across diverse models, highlighting the importance of data generation research as a distinct area from classifier development.
+Tabular data generation is a critical area of research with significant implications [4, 27, 37].
+This task serves two primary purposes: (1) enhancing classification performance in a model-agnostic manner through data augmentation, similar to SMOTE, as shown in Tables 1, 2, and 6, and (2) generating synthetic data to replace original data in security-critical or privacy-sensitive contexts, as demonstrated in Table 7.
+In fields such as healthcare, where collecting new samples or achieving balanced class labels is challenging and data may be noisy or incomplete, generating high-quality synthetic data is crucial.
+Figure 12: Illustration of ablated versions of our proposed method.
+In conclusion, even as more advanced classification models for tabular data are developed in the future, our proposed method could continue to play a crucial role in enhancing performance by generating high-quality synthetic data, thereby potentially making a significant impact on the tabular data classification community and related tasks.
+Appendix C Additional experimental details
+This section provides the illustration of ablated versions (Appendix C.1), comprehensive information on the datasets (Appendix C.2) and comparison baselines (Appendix C.3), implementation details (Appendix C.4), and final prompt design and examples (Appendix C.5).
+C.1 Ablation prompt examples
 This section visualizes five ablated versions of our proposed prompt design, as shown in Fig.
-11.
+12.
 The ablations involve removing or altering four key components: data format, balanced class ratio, multi-class generation, and class grouping.
-Figure 11: Illustration of ablated versions.
-I.2 Dataset
-The detailed description of the datasets are shown in Table 10.
-The variable descriptions are directly sourced from dataset platforms like Kaggle or the UCI repository.
-In our approach, we designate the class with fewer instances as the positive class for binary classification tasks.
-Table 10: Details of the datasets used to evaluate synthetic tabular data generation models.
+Table 12: Dataset details used in this study.
 Dataset #Class #Categorical features #Numerical features #Samples Domain
@@
 Thyroid 2 16 1      364 Medical
-I.3 Baseline reproducibility
+C.2 Dataset
+A detailed description of the datasets is shown in Table 12.
+The variable descriptions are directly sourced from dataset platforms like Kaggle or the UCI repository.
+In our approach, we designate the class with fewer instances as the positive class for binary classification tasks.
+We also provide the class distribution information for the datasets used in this work, as shown in Fig.
+13.
+Among the six datasets, the Travel, Income, Sick, and Thyroid datasets exhibit substantial class imbalances.
+Figure 13: Class distribution for the real-world public tabular datasets used in this study.
+C.3 Baseline reproducibility
 In our research, we utilize implementations from the Synthetic Data Vault [22] for TVAE [35], CTGAN [35], and CopulaGAN, while for the GReaT [4] model, we employ code from its official repository.
-In our experiment, the GReaT model fails to accurately generating column names ‘max_glu_serum’ and ‘A1Cresult’ in the Diabetes dataset, even after training for 85 epochs that reported in the original paper.
+In our experiment, the GReaT model fails to accurately generate column names ‘max_glu_serum’ and ‘A1Cresult’ in the Diabetes dataset, even after training for 85 epochs, as reported in the original paper.
 Extending the training to 105 epochs does not rectify this issue, with ‘max_glu_serum’ and ‘A1Cresult’ being correctly generated in only approximately 2.8% and 12.1% of instances, respectively.
-For these columns, failure cases are treated as NaN value.
+For these columns, failure cases are treated as NaN values.
 We utilize the official open-source code from the official TabDDPM GitHub repository to reproduce CTAB-GAN [39], CTAB-GAN+ [40], and TabDDPM [19].
@@
 Due to the lack of official code provided by the authors of LITO [37], we adopt the prompt design described in their original paper.
-Specifically, we use the prompt for LITO-C to generate minor class data.
+Specifically, we use the prompt for LITO-C to generate minority class data.
 To generate samples for both majority and minority classes, we also use the modified prompt to generate majority class data.
@@
 Although LITO and CuratedLLM propose multi-stage techniques for data generation, we focus solely on evaluating their prompt designs.
-I.4 Experimental setup
+C.4 Experimental setup
 For classification, we utilize the target class of each dataset to divide groups.
-Specifically, the Diabetes dataset has three groups, and the Travel, Sick, HELOC, Income and Thyroid datasets have two groups.
+Specifically, the Diabetes dataset has three groups, and the Travel, Sick, HELOC, Income, and Thyroid datasets have two groups.
 For the process of unique variable mapping, each discrete value is transformed into a combination of three characters, including uppercase letters and digits.
@@
 Where possible, we train the classification models using a single NVIDIA GeForce RTX 3090 GPU on an Ubuntu 18.04 system.
-I.5 More details on the prompt design
-This section provides an in-depth explanation of the construction of the proposed prompt. Prompt examples for generating synthetic data are provided in Tables 11-16.
-Altogether, the final prompt template for synthetic data generation consists of four parts.
+C.5 More details on the prompt design
+This section outlines the construction of the proposed prompt. As shown in the prompt examples in Tables 13-18, the final prompt template for synthetic data generation consists of four parts.
 Descriptions (Optional).
@@
 This header serves as a signal for the LLM to commence generating synthetic data, maintaining the structure and patterns established in the prompt.
-By following this structured approach, the LLM is guided to generate synthetic data that not only adheres to the format and structure of the original dataset but also incorporates variability and distinctiveness in the data patterns.
-The inclusion of descriptions, careful structuring of set-level examples, and strategic repetition of templates all contribute to the effectiveness of this synthetic data generation method.
-Appendix J Confusion matrix results
-This section provides confusion matrix results on four binary classification datasets with Gradient Boosting classifier.
-The confusion matrix in Fig.
-12 reveals that machine learning classification model classifies most accurately when the data generated from our method is added to the original data.
-Figure 12: Confusion matrix on the Travel, Sick, HELOC, and Income datasets using Gradient Boosting classifier.
-Table 11: Example of a data synthesis prompt for the Travel dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data.
+By following this structured approach, the LLM is guided to effectively generate synthetic data that matches the structure and characteristics of the original dataset.
+Appendix D Related work
+LLM in-context learning prompting    With the remarkable advancements in LLMs, extensive research has been conducted in prompt engineering to maximize their potential for various natural language processing (NLP) tasks, including text summarization and question answering, yielding significant value [9, 28, 36, 41].
+Optimizing prompts for specific tasks is inherently a combinatorial challenge, and in the absence of established optimization principles, progress has often been driven by heuristic methods validated through rigorous empirical evaluations [25, 32, 38].
+Our study adopts this empirical approach, conducting extensive experiments to develop effective prompts for synthetic tabular data generation, which holds unique challenges that differ from typical NLP tasks.
+Through these experiments, we introduce a tailored prompt design, EPIC, that employs class-balanced, grouped, and structured formatting and unique variable mapping.
+Tabular data generation    In the field of tabular data synthesis, traditional methods have predominantly focused on interpolation techniques, such as SMOTE [5].
+Although useful, these methods often struggle to capture complex relationships between features [2].
+Recently, generative models have shown promise in generating realistic tabular data [6, 8, 16, 35, 39, 40], and diffusion models have also achieved promising results in this area [19].
+Building on these advances, GReaT [4] achieved notable success by transforming tabular data into natural text format and fine-tuning LLMs for synthetic tabular data generation.
+However, this fine-tuning process is resource-intensive, requiring extensive training of large models for each dataset.
+To address this, LITO [37] and Curated LLM [27] have explored using LLMs through in-context learning for tabular data generation, yet these approaches require additional filtering steps to ensure data quality.
+In contrast, we demonstrate that with comprehensive prompt exploration, high-quality tabular data generation can be achieved through prompt design alone, without the need for additional steps.
+Our work is distinguished by the thorough and comprehensive evaluations across six real-world datasets and synthetic toy data, aiming to provide deeper insights into the prompting process.
+We addressed experiments and analyses often overlooked in prior research on tabular data generation.
+Appendix E Broader Impact
+Our method is generally applicable to tabular data generation and classification tasks with minimal preprocessing, making it accessible for diverse research and industrial applications.
+Potential positive impacts include enhanced data accessibility and privacy preservation, which can support decision-making processes in fields such as healthcare and finance by reducing reliance on real data and protecting individual privacy.
+However, potential negative impacts include the risk of data misuse.
+Synthetic data generation could be exploited maliciously, such as to create deceptive datasets intended to mislead systems or individuals.
+Additionally, using closed-source LLMs poses potential risks, including data leakage through API calls, particularly concerning when handling sensitive data, and limitations on direct access to verify the models being used.
+Ensuring quality, security, and ethical use of generated data under these conditions is essential.
+Appendix F Complete quantitative results
+This section presents the complete confusion matrix results in Appendix F.1 and compares the classification performance with the baselines in Appendix F.2.
+F.1 Confusion matrix results
+We present normalized confusion matrix results on binary classification datasets, as shown in Fig.
+14.
+The results demonstrate that adding the synthetic data generated by our method to the original dataset significantly enhances classification performance, outperforming baselines by a large margin.
+Additionally, using only the synthetic data generated by our method achieves better performance than augmenting the original dataset with synthetic data from baselines.
+F.2 Complete results on comparison with baselines
+Table 19 offers the complete performance comparison results with standard deviation values of 20 independent experiments on all six datasets.
+Figure 14: Normalized confusion matrix results on binary classification datasets using the gradient boosting classifier. Ours only denotes cases where only our synthetic data are used.
+Table 13: Example of an EPIC prompt for the Travel dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data.
 Template Prompt sample Descriptions Churn: whether customer churns or doesnt churn for tour and travels company,
@@
 Set Header (trigger) Churn,Age,FrequentFlyer,AnnualIncomeClass,ServicesOpted, AccountSyncedToSocialMedia,BookedHotelOrNot
-Table 12: Example of a data synthesis prompt for the Sick dataset.
+Table 14: Example of an EPIC prompt for the Sick dataset.
 This prompt illustrates the structure of sets and groups.
@@
 Set Header (trigger) Class,age,sex,on_thyroxine,query_on_thyroxine, on_antithyroid_medication,sick,pregnant,thyroid_surgery, I131_treatment,query_hypothyroid,query_hyperthyroid,lithium, goitre,tumor,hypopituitary,psych,TSH_measured,TSH,T3_measured,T3, TT4_measured,TT4,T4U_measured,T4U,FTI_measured,FTI,referral_source
-Table 13: Example of a data synthesis prompt for the HELOC dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data.
+Table 15: Example of an EPIC prompt for the HELOC dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data.
 Template Prompt sample Descriptions Set Header RiskPerformance,ExternalRiskEstimate,MSinceOldestTradeOpen, MSinceMostRecentTradeOpen,AverageMInFile,NumSatisfactoryTrades, NumTrades60Ever2DerogPubRec,NumTrades90Ever2DerogPubRec, PercentTradesNeverDelq,MSinceMostRecentDelq, MaxDelq2PublicRecLast12M,MaxDelqEver,NumTotalTrades, NumTradesOpeninLast12M,PercentInstallTrades, MSinceMostRecentInqexcl7days,NumInqLast6M,NumInqLast6Mexcl7days, NetFractionRevolvingBurden,NetFractionInstallBurden, NumRevolvingTradesWBalance,NumInstallTradesWBalance, NumBank2NatlTradesWHighUtilization,PercentTradesWBalance Group A.
@@
 Set Header (trigger) RiskPerformance,ExternalRiskEstimate,MSinceOldestTradeOpen, MSinceMostRecentTradeOpen,AverageMInFile,NumSatisfactoryTrades, NumTrades60Ever2DerogPubRec,NumTrades90Ever2DerogPubRec, PercentTradesNeverDelq,MSinceMostRecentDelq, MaxDelq2PublicRecLast12M,MaxDelqEver,NumTotalTrades, NumTradesOpeninLast12M,PercentInstallTrades, MSinceMostRecentInqexcl7days,NumInqLast6M,NumInqLast6Mexcl7days, NetFractionRevolvingBurden,NetFractionInstallBurden, NumRevolvingTradesWBalance,NumInstallTradesWBalance, NumBank2NatlTradesWHighUtilization,PercentTradesWBalance
-Table 14: Example of a data synthesis prompt for the Income dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data.
+Table 16: Example of an EPIC prompt for the Income dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data.
 Template Prompt sample Descriptions Set Header income,age,workclass,fnlwgt,education,education_num, marital-status,occupation,relationship,race,sex,capital-gain, capital-loss,hours-per-week,native-country Group A.
@@
 Set Header (trigger) income,age,workclass,fnlwgt,education,education_num, marital-status,occupation,relationship,race,sex,capital-gain, capital-loss,hours-per-week,native-country
-Table 15: Example of a data synthesis prompt for the Diabetes dataset.
+Table 17: Example of an EPIC prompt for the Diabetes dataset.
 This prompt illustrates the structure of sets and groups.
@@
 […]
-Table 16: Example of a data synthesis prompt for the Thyroid dataset.
-This prompt illustrates the structure of sets and groups.
-Each iteration utilizes randomly selected samples from the training data.
+Table 18: Example of an EPIC prompt for the Thyroid dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data.
 Template Prompt sample Descriptions Set Header Recurred,Age,Gender,Smoking,Hx Smoking,Hx Radiothreapy, Thyroid Function,Physical Examination,Adenopathy,Pathology, Focality,Risk,T,N,M,Stage,Response Group A.
@@
 Set Header (trigger) Recurred,Age,Gender,Smoking,Hx Smoking,Hx Radiothreapy, Thyroid Function,Physical Examination,Adenopathy,Pathology, Focality,Risk,T,N,M,Stage,Response
-Appendix K Complete results on comparison with baseline methods
-Table K offers the complete performance comparison results with standard deviation values of 20 independent experiments on all six datasets.
-Table 17: Complete results on comparison of classification performance improvement using synthetic data from baseline models and our method, averaged across four classifiers: XGBoost, CatBoost, LightGBM, and Gradient Boosting classifier, with each model run five times.
-“#syn” represents the number of synthetic samples added to the original data.
+Table 19: Complete results comparing ML classification performance when synthetic data are added to the original dataset.
+Results are averaged across four classifiers: XGBoost, CatBoost, LightGBM, and the gradient boosting classifier, with each model run five times.
+#syn denotes the number of synthetic samples added to the original dataset.
 Dataset Method #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
@@
  +Ours +1K  94.80$\pm$1.02 (+0.57) 96.39$\pm$0.62 (+1.31) 95.23$\pm$1.02 (+4.09) 97.55$\pm$0.87 (-1.47)
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The Abstract and Introduction (Section 1) accurately reflect the contributions and scope of our paper. Refer to the contribution summary at the end of the Introduction.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: The limitations are provided in Section 4.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: This work does not include theoretical results.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer:[Yes]
+  19. Justification: The paper includes comprehensive experimental details in Section 3 and Appendix C, and provides supplementary code to facilitate reproducibility.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We publicly release the source code, along with instructions to reproduce the experimental results.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: The paper specifies the training and test details, including data splits, hyperparameters, and their selection process, in Section 3 and Appendix C.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: The paper reports standard deviations in Fig. 2 and ML classification performance tables in main manuscript, and the experimental results in Section A and Section F.2.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We provide information about the GPU resources used for the experiments in Appendix C.4.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We have reviewed and adhered to the NeurIPS Code of Ethics throughout the research process.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We discuss broader impacts in Appendix E.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: No data or models are being released in this paper, so there is no risk for misuse.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: All existing assets used in the paper are properly credited.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: The source code includes a README file that provides comprehensive documentation.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.